### v1

# Max-Distance Sparsification for Diversification and Clustering

Abstract
Let $\mathcal{D}$ be a set family that is the solution domain of some combinatorial problem. The max-min diversification problem on $\mathcal{D}$ is the problem to select $k$ sets from $\mathcal{D}$ such that the Hamming distance between any two selected sets is at least $d$. FPT algorithms parameterized by $k,l:=\max_{D\in\mathcal{D}}|D|$ and $k,d$ have been actively studied recently for several specific domains.
This paper provides unified algorithmic frameworks to solve this problem. Specifically, for each parameterization $k,l$ and $k,d$, we provide an FPT oracle algorithm for the max-min diversification problem using oracles related to $\mathcal{D}$. We then demonstrate that our frameworks generalize most of the existing domain-specific tractability results and provide the first FPT algorithms for several domains.
Our main technical breakthrough is introducing the notion of max-distance sparsifier of $\mathcal{D}$, a domain on which the max-min diversification problem is equivalent to the same problem on the original domain $\mathcal{D}$. The core of our framework is to design FPT oracle algorithms that construct a constant-size max-distance sparsifier of $\mathcal{D}$. Using max-distance sparsifiers, we provide FPT algorithms for the max-min and max-sum diversification problems on $\mathcal{D}$, as well as $k$-center and $k$-sum-of-radii clustering problems on $\mathcal{D}$, which are also natural problems in the context of diversification and have their own interests.
1 Introduction
1.1 Background and Motivation
The procedure for approaching real-world problems with optimization algorithms involves formulating the real-world motivations as mathematical problems and then solving them. However, real-world problems are complex, and the idea of a “good” solution cannot always be correctly formulated. Attempting to impose a formulation may ignore important perspectives such as human sensibilities and ethical considerations. The paradigm of diversification, introduced by [baste2019fpt] and [baste2022diversity], is a “formulation of the unformulatable problems”, which formulates diversity measures for a set of multiple solutions, rather than attempting to formulate the “goodness” of a single solution. By computing a set of solutions that maximize this measure, the algorithm provides effective options to evaluators who have the correct criteria for judging the “goodness” of a solution.
Let $U$ be a finite set, $k\in\mathbb{Z}_{\geq 1}$ and $d\in\mathbb{Z}_{\geq 0}$. Let $\mathcal{D}\subseteq 2^{U}$ be the feasible domain of some combinatorial problem. The following problem frameworks, defined by two types of diversity measures, have been studied extensively.
Max-Min Diversification Problem on $\mathcal{D}$: Does there exist a $k$-tuple $(D_{1},\dots,D_{k})\in\mathcal{D}^{k}$ of sets in $\mathcal{D}$ such that $\min_{1\leq i<j\leq k}|D_{i}\triangle D_{j}|\geq d$? [^1] We define $Z_{1}\triangle Z_{2}:=(Z_{1}\setminus Z_{2})\cup(Z_{2}\setminus Z_{1})$.
Max-Sum Diversification Problem on $\mathcal{D}$: Does there exist a $k$-tuple $(D_{1},\dots,D_{k})\in\mathcal{D}^{k}$ of sets in $\mathcal{D}$ such that $\sum_{1\leq i<j\leq k}|D_{i}\triangle D_{j}|\geq d$?
These problems ensure diversity by aiming to output solutions that are as dissimilar as possible in terms of Hamming distance.
Parameterized algorithms for diversification problems have been actively studied. Particularly, FPT algorithms for the max-min diversification problems parameterized by $k$ and $l:=\max_{D\in\mathcal{D}}|D|$ [baste2022diversity, baste2019fpt, fomin2024diversecollection, hanaka2021finding, shida2024finding], as well as by $k$ and $d$ [eiben2024determinantal, fomin2024diversepair, fomin2024diversecollection, funayama2024parameterized], have been the focus of research. Since assuming $d\leq 2l$ does not lose generality in the max-min diversification problem, the latter addresses a more general situation than the former. The max-sum diversification problem is empirically more tractable than the max-min diversification problem, and most of the existing algorithms for the max-min diversification problems can be straightforwardly converted to solve the max-sum diversification problems on the corresponding domain. Therefore, for some time hereafter, we will restrict our discussion to the max-min diversification problem.
The most successful technical framework in this field is color-coding [alon1995color]. [hanaka2021finding] constructed a color-coding based framework designing FPT algorithms parameterized by $k,l$ for the max-min diversification problem on several domains. [shida2024finding] provided the same result for the case where $\mathcal{D}$ is the family of longest common subsequences. Color-coding has also been successful for the case where the parameter is $k,d$. [fomin2024diversepair] provided a color-coding based FPT algorithm for the max-min diversification problem on $\mathcal{D}$ parameterized by $d$, where $\mathcal{D}$ is the family of maximum matchings and $k=2$. Furthermore, [fomin2024diversecollection] provided a color-coding based FPT algorithm parameterized by $k,d$ for the family of perfect matchings, and [funayama2024parameterized] provided the same result for the family of shortest paths. All of these algorithms use a subroutine to find a colorful set that belongs to a domain related to $\mathcal{D}$ for a given coloring of $U$.
This research provides general algorithmic frameworks for FPT algorithms addressing max-min (and max-sum) diversification problems for both parameterizations $k,l$ and $k,d$. Our framework is very general and can be applied to all domains [bang2021k, bang2016parameterized, baste2022diversity, baste2019fpt, eiben2024determinantal, fomin2024diversepair, fomin2024diversecollection, funayama2024parameterized, gutin2018k, hanaka2021finding, shida2024finding] for which FPT algorithms parameterized by $k,l$ and $k,d$ are currently known for the case that diversity measure is defined using an unweighted Hamming distance and several domains where such algorithms were previously unknown.
Unlike existing studies, our framework is not based on color-coding. Very roughly speaking, our framework solves the max-min diversification problem on $\mathcal{D}$ by utilizing an oracle that determines whether there exists a set of a specified size in the domain related to $\mathcal{D}$ that satisfies additional simple constraints. The design of this oracle is generally much simpler than the design of an oracle searching for colorful sets in the same domain, making our framework easy-to-use and applicable for a broader domain $\mathcal{D}$.
Our main technical breakthrough is introducing a notion of max-distance sparsifier as an intermediate step for solving the max-min diversification problem, whose precise definition is given in Section 1.3. The most critical fact is that when $\mathcal{K}$ is a max-distance sparsifier of $\mathcal{D}$, the max-min diversification problem on $\mathcal{D}$ is equivalent to the same problem on $\mathcal{K}$. Our framework constructs a max-distance sparsifier $\mathcal{K}$ of $\mathcal{D}$ with a constant size, enabling us to solve the max-min diversification problems on $\mathcal{D}$ by brute-force search on $\mathcal{K}$.
The power of max-distance sparsification is not limited to solving max-min (and max-sum) diversification problems. Specifically, the following $k$-center [hsu1979easy] and $k$-sum-of-radii clustering problems on $\mathcal{D}$ [charikar2001clustering] can also be solved via max-distance sparsification.
$k$-Center Clustering Problem on $\mathcal{D}$: Does there exist a $k$-tuple of subsets $(D_{1},\dots,D_{k})\in\mathcal{D}^{k}$ such that for all $D\in\mathcal{D}$, there exists an $i\in\{1,\dots,k\}$ satisfying $|D_{i}\triangle D|\leq d$?
$k$-Sum-of-Radii Clustering Problem on $\mathcal{D}$: Does there exist a $k$-tuple of subsets $(D_{1},\dots,D_{k})\in\mathcal{D}^{k}$ and a $k$-tuple of non-negative integers $(d_{1},\dots,d_{k})\in\mathbb{Z}_{\geq 0}^{k}$ with $\sum_{i\in\{1,\dots,k\}}d_{i}\leq d$ such that for all $D\in\mathcal{D}$, there exists an $i\in\{1,\dots,k\}$ satisfying $|D_{i}\triangle D|\leq d_{i}$?
When $\mathcal{D}$ is an explicitly given set of points, parameterized algorithms for these problems have been extensively studied in the area of clustering [amir2014efficiency, bandyapadhyayL023a, chen2024parameterized, demaine2005fixed, eiben2023parameterized, feldmann2020parameterized, inamdar2020capacitated]. This research initiates the study of clustering problems when $\mathcal{D}$ is implicitly given as the solution domain of combinatorial problems. These problems are also natural in the context of diversification, since in real situations, the concept of diversity often means that the extracted elements cover the entire space comprehensively rather than being mutually dissimilar. This motivation is formulated by clustering problems, which extract a list of sets in $\mathcal{D}$ such that for each set in $\mathcal{D}$, there is an extracted set near to it.
1.2 Our Results
This paper consists of two parts. In the first part, we design general frameworks for solving diversification and clustering problems. In the second part, we apply our frameworks to several specific domains $\mathcal{D}$.
1.2.1 The Frameworks
We define the following $(-1,1)$-optimization oracle on $\mathcal{D}$ and the exact extension oracle on $\mathcal{D}$.
$(-1,1)$-Optimization Oracle on $\mathcal{D}$: Let $U$ be a finite set, $\mathcal{D}\subseteq 2^{U}$, and $w\in\{-1,1\}^{U}$ be a weight vector. Return a set $D\in\mathcal{D}$ that maximizes $\sum_{e\in D}w_{e}$.
Exact Extension Oracle on $\mathcal{D}$: Let $U$ be a finite set, $r\in\mathbb{Z}_{\geq 0}$, $\mathcal{D}\subseteq 2^{U}$, and $C\in\mathcal{D}$. Let $X,Y\subseteq U$ be two disjoint subsets of $U$. If there exists a set $D\in\mathcal{D}$ such that $|D\triangle C|=r$, $X\subseteq D$, and $Y\cap D=\emptyset$, return one such set. If no such set exists, return $\bot$.
When $C=\emptyset$ and $X=\emptyset$, we specifically call the exact extension oracle on $\mathcal{D}$ the exact empty extension oracle on $\mathcal{D}$.
Exact Empty Extension Oracle on $\mathcal{D}$: Let $U$ be a finite set, $r\in\mathbb{Z}_{\geq 0}$, $\mathcal{D}\subseteq 2^{U}$, and $Y\subseteq U$. If there exists a set $D\in\mathcal{D}$ such that $|D|=r$ and $Y\cap D=\emptyset$, return one such set. If no such set exists, return $\bot$.
Let $\mathcal{P}_{\mathcal{D}}$ be the max-min/max-sum diversification problem or the $k$-center/$k$-sum-of-radii clustering problem on $\mathcal{D}$. Our main result is FPT algorithms for solving $\mathcal{P}_{\mathcal{D}}$ using these oracles. We construct frameworks for both types of parameterizations, $k,l$ and $k,d$. The result for the parameterization by $k,l$ is as follows.
Theorem 1.1.
There exists an oracle algorithm solving $\mathcal{P}_{\mathcal{D}}$ using the exact empty extension oracle on $\mathcal{D}$, where the number of oracle calls and time complexity are both FPT parameterized by $k,l$ and for each call of an oracle, $r$ and $|Y|$ are bounded by constants that depend only on $k,l$.
The result for the parameterization by $k,d$ is as follows.
Theorem 1.2.
There exists a randomized oracle algorithm solving $\mathcal{P}_{\mathcal{D}}$ using the $(-1,1)$-optimization oracle on $\mathcal{D}$ and the exact extension oracle on $\mathcal{D}$, where the number of oracle calls and time complexity are both FPT parameterized by $k,d$ and for each call of the exact extension oracle, $r,|X|,|Y|$ are bounded by constants that depend only on $k,d$.
1.2.2 Applications of Theorem 1.1
For most domains $\mathcal{D}$, the exact empty extension oracle on $\mathcal{D}$ can be straightforwardly designed by using an algorithm for extracting a single element from $\mathcal{D}$. Combining Theorem 1.1 with this empirical fact, we can claim the following surprising fact: For most domain $\mathcal{D}$, the diversification and clustering problems parameterized by $k,l$ on $\mathcal{D}$ are as easy as determining the non-emptiness of $\mathcal{D}$. We demonstrate this by providing an FPT algorithm for the exact empty extension oracle on several domains $\mathcal{D}$ on which an FPT algorithm for the max-min diversification problem parameterized by $k,l$ is known. Specifically, we design FPT algorithms for $\mathcal{D}$ being the vertex cover [baste2022diversity, baste2019fpt], $t$-hitting set [baste2022diversity], feedback vertex set [baste2022diversity], and common independent set of two matroids [fomin2024diversecollection]. Furthermore, we prove that our framework generalizes existing frameworks provided by [baste2019fpt] and [hanaka2021finding] in both the broader applicability to more general domains $\mathcal{D}$ and the applicability to clustering problems.
1.2.3 Applications of Theorem 1.2
FPT algorithms for the max-min diversification problems on $\mathcal{D}$ parameterized by $k,d$ are known for the cases where $\mathcal{D}$ is the family of matroid bases [fomin2024diversecollection], perfect matchings [fomin2024diversecollection], and shortest paths [funayama2024parameterized]. Additionally, FPT algorithms parameterized by $k,l$ are known for the cases where $\mathcal{D}$ is the family of interval schedulings [hanaka2021finding] and the longest common subsequences of absolute constant number of strings [shida2024finding]. We consider the following domains $\mathcal{D}$, which include, generalize, or are entirely novel compared to, the cases mentioned above.
Matroid Base Domain: Let $M=(U,\mathcal{I})$ be a matroid. $\mathcal{D}\subseteq 2^{U}$ is the family of bases of $M$.
Branching Domain: Let $G=(V,E)$ be a directed graph and $r\in V$. $\mathcal{D}\subseteq 2^{E}$ is the family of edge sets of branchings with root $r$ in $G$.
Matching Domain: Let $G=(V,E)$ be an undirected graph and $l\in\mathbb{Z}_{\geq 0}$. $\mathcal{D}\subseteq 2^{E}$ is the family of matchings of size $l$ in $G$.
Minimum Edge $s,t$-Flow Domain: Let $G=(V,E)$ be a directed graph, $b\in\mathbb{Z}_{\geq 0}$, and $s,t\in V$. $\mathcal{D}\subseteq 2^{E}$ is the family of the edge sets of $s,t$-flows of amount $b$ using minimum number of edges.
Minimum Steiner Tree Domain: Let $G=(V,E)$ be an undirected graph and $T\subseteq V$. $\mathcal{D}\subseteq 2^{E}$ is the family of edge sets of minimum Steiner trees in $G$ that connects the terminal set $T$.
Minimum $s,t$-Cut Domain: Let $G=(V,E)$ be a directed graph and $s,t\in V$. $\mathcal{D}\subseteq 2^{V}$ is the family of vertex sets $Z\subseteq V$ that include $s$, exclude $t$, and minimize the number of edges between $Z$ and $V\setminus Z$.
Edge Bipartization Domain: Let $G=(V,E)$ be an undirected graph. $\mathcal{D}\subseteq 2^{V}$ is the family of vertex sets $Z\subseteq V$ such that the total number of edges either within $Z$ or within $V\setminus Z$ takes the minimum value $s$.
Dynamic Programming Problems Domain: Let $U$ be a finite set, $G=(V,E)$ be a directed acyclic graph, and $q\in V^{U}$ be a labeling of vertices such that no path in $G$ passes through multiple vertices with the same label. $\mathcal{D}\subseteq 2^{U}$ is the family of subsets $D\subseteq U$ such that there exists a longest path in $G$ where the set of labels of the vertices on the path is $D$.
See Section 5.2 for the more precise definitions of each domain.
By applying Theorem 1.2, we provide FPT algorithms for diversification and clustering problems on all those domains $\mathcal{D}$, where the parameterization is $k$ and $d$ except for the minimum Steiner tree domain, which is parameterized by $k,d,|T|$, and the edge bipartization domain, which is parameterized by $k,d,s$. This result generalizes several existing results in both the broader applicability to more general domain $\mathcal{D}$ and the applicability to clustering problems. Specifically, for the matching domain, our result for the case of $|V|=2l$ corresponds to the result in [fomin2024diversecollection] for perfect matchings, and for the minimum edge $s,t$-flow domain and minimum Steiner tree domain, our results for the cases of $b=1$ and $|T|=2$, respectively, correspond to the results in [funayama2024parameterized] for the shortest paths. Our result for the dynamic programming problems domain implies that for a wide range of problems solvable by dynamic programming, the corresponding diversification and clustering problems have FPT algorithms parameterized by $k$ and $d$. Particularly, this improves the parameterization of the result in [hanaka2021finding] for the interval schedulings and the result in [shida2024finding] for the longest common subsequences from $k,l$ to $k,d$. Furthermore, the technique for the branching domain leads to another FPT algorithm for the $d$-distinct branchings problem [bang2021k, bang2016parameterized, gutin2018k], which asks to find a pair consisting of a branching and an in-branching such that the Hamming distance between them is at least $d$. We also generalize this algorithm to the version of choosing multiple branchings and in-branchings, which is a new result in the literature of the $d$-distinct branchings problem. The rest of our results are entirely new, even for the max-min diversification problem.
1.3 Framework Overview
In this section, we provide an overview of the entire flow of our frameworks.
1.3.1 From $d$-Limited $k$-Max-Distance Sparsifier to Diversification and Clustering
The key idea for Theorem 1.1 is to design the following $k$-max-distance sparsifier of $\mathcal{D}$.
Definition 1.3 ($k$-max-distance sparsifier).
Let $k\in\mathbb{Z}_{\geq 1}$. Let $U$ be a finite set and $\mathcal{D},\mathcal{F}\subseteq 2^{U}$. We say that $\mathcal{K}\subseteq\mathcal{D}$ is a $k$-max-distance sparsifier of $\mathcal{D}$ with respect to $\mathcal{F}$ if for any $(F_{1},\dots,F_{k})\in\mathcal{F}^{k}$ and $(z_{1},\dots,z_{k})\in\mathbb{Z}_{\geq 0}^{k}$, the two conditions
  • There exists $D\in\mathcal{D}$ such that for each $i\in\{1,\dots,k\}$, $|F_{i}\triangle D|\geq z_{i}$.
  • There exists $K\in\mathcal{K}$ such that for each $i\in\{1,\dots,k\}$, $|F_{i}\triangle K|\geq z_{i}$.
are equivalent. Unless specifically noted, when we write $k$-max-distance sparsifier of $\mathcal{D}$, we mean the case where $\mathcal{D}=\mathcal{F}$.
Similarly, the key idea for Theorem 1.2 is to design the following $d$-limited $k$-max-distance sparsifier of $\mathcal{D}$.
Definition 1.4 ($d$-limited $k$-max-distance sparsifier).
Let $k\in\mathbb{Z}_{\geq 1}$ and $d\in\mathbb{Z}_{\geq 0}$. Let $U$ be a finite set and $\mathcal{D},\mathcal{F}\subseteq 2^{U}$. We say that $\mathcal{K}\subseteq\mathcal{D}$ is a $d$-limited $k$-max-distance sparsifier of $\mathcal{D}$ with respect to $\mathcal{F}$ if for any $(F_{1},\dots,F_{k})\in\mathcal{F}^{k}$ and $(z_{1},\dots,z_{k})\in\{0,\dots,d\}^{k}$, the two conditions
  • There exists $D\in\mathcal{D}$ such that for each $i\in\{1,\dots,k\}$, $|F_{i}\triangle D|\geq z_{i}$.
  • There exists $K\in\mathcal{K}$ such that for each $i\in\{1,\dots,k\}$, $|F_{i}\triangle K|\geq z_{i}$.
are equivalent. Unless specifically noted, when we write $d$-limited $k$-max-distance sparsifier of $\mathcal{D}$, we mean the case where $\mathcal{D}=\mathcal{F}$.
The difference between the two sparsifiers is that the domain of $(z_{1},\dots,z_{k})$ is $\mathbb{Z}_{\geq 0}^{k}$ in the former case, while it is $\{0,\dots,d\}^{k}$ in the latter. By definition, any $k$-max-distance sparsifier is also a $d$-limited $k$-max-distance sparsifier for any $d\in\mathbb{Z}_{\geq 0}$.
We can prove that given a $d$-limited $(k-1)$-max-distance sparsifier of $\mathcal{D}$ with size bounded by a constant that depends only on $k$ and $d$, we can construct FPT algorithms parameterized by $k$ and $d$ for the max-min/max-sum diversification problems on $\mathcal{D}$. Similarly, we can prove that given a $(d+1)$-limited $k$-max-distance sparsifier of $\mathcal{D}$ with size bounded by a constant that depends only on $k$ and $d$, we can construct FPT algorithms parameterized by $k$ and $d$ for the $k$-center/$k$-sum-of-radii clustering problems on $\mathcal{D}$. Therefore, to prove Theorems 1.1 and 1.2, it suffices to construct FPT oracle algorithms for designing $k$-max-distance sparsifiers and $d$-limited $k$-max-distance sparsifiers, respectively.
1.3.2 Computing $k$-Max-Distance Sparsifier
The remaining task towards Theorem 1.1 is to provide an FPT algorithm parameterized by $k,l$ that constructs a $k$-max-distance sparsifier of $\mathcal{D}$ with size bounded by a constant that depends only on $k$ and $l$. Our algorithm starts with $\mathcal{K}=\emptyset$ and exhaustively adds sets of $\mathcal{D}$ to $\mathcal{K}$ until $\mathcal{K}$ becomes a $k$-max-distance sparsifier.
We can prove that if adding a set to $\mathcal{K}$ would cause a sufficiently large sunflower (see Section 3 for the definition) consisting of sets of the same size to appear in $\mathcal{K}$, then we do not need to add that set to $\mathcal{K}$ (in fact, for the sake of simplifying the framework, we prove a slightly stronger statement). By using the well-known sunflower lemma (Lemma 3.1), we can bound the number of repetitions by a constant. To choose a set to be added, we design an FPT algorithm using calls of the exact empty extension oracle.
1.3.3 Computing $d$-Limited $k$-Max-Distance Sparsifier
The algorithm in the previous section alone is insufficient to prove Theorem 1.2 since $l$ is unbounded and the bound for the number of repetitions cannot be used. Our algorithm divides $\mathcal{D}$ into at most $k$ clusters, computes a $d$-limited $k$-max-distance sparsifier for each cluster, and outputs their union. Let $p>2d$ be a constant that depends only on $k$ and $d$. We first find $\mathcal{C}\subseteq\mathcal{D}$ satisfying the following properties.
  • $|\mathcal{C}|\leq k$.
  • For all distinct $C,C^{\prime}\in\mathcal{C}$, $|C\triangle C^{\prime}|>2d$.
  • For all $D\in\mathcal{D}$, there exists $C\in\mathcal{C}$ such that $|D\triangle C|\leq p$.
If such a family does not exist, a trivial $d$-limited $k$-max-distance sparsifier of $\mathcal{D}$ will be found, and we output it and terminate.
We provide an algorithm for computing $\mathcal{C}$. Our algorithm starts with $\mathcal{C}=\emptyset$ and exhaustively adds sets in $\mathcal{D}$ to $\mathcal{C}$ until $\mathcal{C}$ satisfies the above conditions or its size exceeds $k$. To choose the elements to be added, we randomly sample $w\in\{-1,1\}^{U}$ and call a $(-1,1)$-optimization oracle. We can prove for sufficiently large constant $p$ that if there exists $D\in\mathcal{D}$ such that $|D\triangle C|>p$ for any $C\in\mathcal{C}$, with a constant probability, the $(-1,1)$-optimization oracle will find a $D\in\mathcal{D}$ such that $|D\triangle C|>2d$ for any $C\in\mathcal{C}$. Thus, if $\mathcal{C}$ does not meet the conditions, by calling the $(-1,1)$-optimization oracle a sufficient number of times, we can find a set to add to $\mathcal{C}$ with high probability.
Here, we provide an algorithm for computing a $d$-limited $k$-max-distance sparsifier of $\mathcal{D}$ using $\mathcal{C}$. For each cluster $\mathcal{D}_{C}:=\{D\in\mathcal{D}\colon|D\triangle C|\leq p\}$, let $\mathcal{D}^{*}_{C}:=\{D\triangle C\colon D\in\mathcal{D}_{C}\}$. The algorithm computes a $k$-max-distance sparsifier of each $\mathcal{D}^{*}_{C}$ and outputs their union. For technical reasons, we actually compute a slightly more general object, but we will not delve into the details here. Since each $\mathcal{D}^{*}_{C}$ consists only of sets whose size is at most $p$, the $k$-max-distance sparsifier of $\mathcal{D}^{*}_{C}$ can be constructed using the algorithm in Section 1.3.2. The exact empty extension oracle on $\mathcal{D}^{*}_{C}$ corresponds to the exact extension oracle on $\mathcal{D}$.
Here, we note the difference between our framework and that used by [fomin2024diversecollection] and [funayama2024parameterized] to provide FPT algorithms for the max-min diversification problem on $\mathcal{D}$ when $\mathcal{D}$ is the family of perfect matchings and shortest paths, respectively. Their algorithms also start by dividing $\mathcal{D}$ into clusters. However, their algorithms perform stricter clustering than ours. Specifically, in their clustering, clusters $\mathcal{D}_{C}$ corresponding to different $C\in\mathcal{C}$ must be well-separated. In contrast, we allow clusters to overlap. This simplifies the clustering step compared to their approach at the cost of a more challenging task afterward. We resolve this more challenging task by introducing and designing the $d$-limited $k$-max-distance sparsifier.
1.4 Further Related Work
1.4.1 Max-Sum Diversification Problems
The max-sum diversification problem is often more tractable compared to the max-min diversification problem, and polynomial-time algorithms are known for multiple domains $\mathcal{D}$. [hanaka2021finding] provided a polynomial-time algorithm for the case where $\mathcal{D}$ is the base family of a matroid. [hanaka2022computing] provided polynomial-time algorithms for the cases where $\mathcal{D}$ is the family of shortest paths, branchings, and bipartite matchings. [BergMS23] provided a polynomial-time algorithm for the case where $\mathcal{D}$ is the family of edge sets of minimum $s,t$-cuts.
There has also been active research on approximation algorithms for the max-sum diversification problem. [hanaka2023framework] proposed a generic framework that provides local search-based approximation algorithms for max-sum diversification problems. [gao2022obtaining] provided a framework for bicriteria approximation algorithms for the case where $\mathcal{D}$ is a family of (not necessarily optimal) solutions of an optimization problem. [do2023diverse] discussed the tradeoff between solution quality and diversity for submodular maximization on matroids.
1.4.2 Further Parameterized Algorithms for Max-Min Diversification Problems
There are several other research directions on parameterized algorithms for max-min diversification problems. [eiben2024determinantal] introduced a technique called determinantal sieving and improved the time complexity of the algorithms by [fomin2024diversecollection] for the cases where $\mathcal{D}$ is the base of a matroid, a matching, and the common independent set of two matroids. [drabik2024finding] provided an FPT algorithm for diversification problems on domains expressible by $\mathrm{MSO}_{1}$ formulas on graphs with bounded cliquewidth, parameterized by $k$, cliquewidth, and the length of the $\mathrm{MSO}_{1}$ formula. [arrighi2023synchronization] proposed an FPT algorithm for the max-min diversification problem on the set of synchronizing words for a deterministic finite automaton, where the distance between two words is measured by the edit distance.
1.4.3 Parameterized Algorithms for Clustering Problems
For the case where the point set is explicitly given, several researches have been conducted on parameterized algorithms for clustering problems. It is easy to see that when parameterized solely by $k$ and the metric is general, the $k$-center clustering problem includes the dominating set problem, which is known to be W[2]-hard. Therefore, the study of FPT approximation algorithms is very active recently [bandyapadhyayL023a, chen2024parameterized, eiben2023parameterized, feldmann2020parameterized, inamdar2020capacitated]. Research on providing exact parameterized algorithms on restricted metrics is also being conducted [demaine2005fixed]. Probably, the research most closely related to our situation is the extension of the closest string problem by [amir2014efficiency]. They defined the $k$-center and $k$-sum-of-radii clustering problems for sets of strings and investigated their parameterized complexity. When the alphabet is binary, they consider the same problem as ours. However, their results do not directly imply our results since, in our setting, the points are implicitly given as the solution domain of combinatorial problems, which are generally of exponential size.
1.5 Organization
The rest of this paper is organized as follows. In Section 2, we provide FPT algorithms for solving diversification and clustering problems on $\mathcal{D}$ using a $d$-limited $k$-max-distance sparsifier of $\mathcal{D}$ of constant size. In Section 3, we complete the proof of Theorem 1.1 by providing an FPT oracle algorithm parameterized by $k,l$ that computes a $k$-max-distance sparsifier of $\mathcal{D}$ with size bounded by a constant that depends only on $k,l$. In Section 4, we complete the proof of Theorem 1.2 by providing an FPT oracle algorithm parameterized by $k,d$ that computes a $d$-limited $k$-max-distance sparsifier of $\mathcal{D}$ with size bounded by a constant that depends only on $k,d$. The discussion in Section 4 internally uses the results from Section 3. Finally, in Section 5, we apply the results of Theorems 1.1 and 1.2 to several domains $\mathcal{D}$ to obtain FPT algorithms for diversification and clustering problems.
2 From Sparsifier to Diversification and Clustering
In this section, we provide FPT algorithms for diversification and clustering problems using a $d$-limited $k$-max-distance sparsifier of constant size.
2.1 Diversification
Let $U$ be a finite set, $d\in\mathbb{Z}_{\geq 0}$, $k\in\mathbb{Z}_{\geq 1}$, and $\mathcal{D}\subseteq 2^{U}$. For diversification problems, we have the following.
Lemma 2.1.
Let $\mathcal{K}\subseteq\mathcal{D}$ be a $d$-limited $(k-1)$-max-distance sparsifier of $\mathcal{D}$ and $(D_{1},\dots,D_{k})\in\mathcal{D}^{k}$. Then, there is a $k$-tuple $(K_{1},\dots,K_{k})\in\mathcal{K}^{k}$ such that $\min(d,|D_{i}\triangle D_{j}|)\leq\min(d,|K_{i}\triangle K_{j}|)$ holds for all $1\leq i<j\leq k$.
Proof.
Assume $(D_{1},\dots,D_{k})\in\mathcal{D}^{k}\setminus\mathcal{K}^{k}$ and let $i\in\{1,\dots,k\}$ be an index such that $D_{i}\not\in\mathcal{K}$. It is sufficient to prove that there is a set $K_{i}\in\mathcal{K}$ such that $\min(d,|D_{i}\triangle D_{j}|)\leq\min(d,|K_{i}\triangle D_{j}|)$ holds for all $j\in\{1,\dots,k\}\setminus\{i\}$. For $j\in\{1,\dots,k\}\setminus\{i\}$, let $z_{j}:=\min(d,|D_{i}\triangle D_{j}|)$. Since $\mathcal{K}$ is a $d$-limited $(k-1)$-max-distance sparsifier of $\mathcal{D}$, there exists $K_{i}\in\mathcal{K}$ such that $\min(d,|K_{i}\triangle D_{j}|)\geq\min(d,z_{j})=\min(d,|D_{i}\triangle D_{j}|)$ holds for all $j\in\{1,\dots,k\}\setminus\{i\}$. ∎
Therefore, we have the following.
Lemma 2.2.
Let $\mathcal{K}\subseteq\mathcal{D}$ be a $d$-limited $(k-1)$-max-distance sparsifier of $\mathcal{D}$. If there exists a solution to the max-min/max-sum diversification problem on $\mathcal{D}$, then there exists a solution consisting only of sets in $\mathcal{K}$.
Considering an algorithm that exhaustively searches for a subfamily of $\mathcal{K}$ of size $k$, we can state the following.
Lemma 2.3.
Assume there exists an FPT algorithm parameterized by $d,k$ to compute a $d$-limited $(k-1)$-max-distance sparsifier of $\mathcal{D}$ with size bounded by a constant that depends only on $d,k$. Then, there exists an FPT algorithm parameterized by $d,k$ for the max-min/max-sum diversification problem on $\mathcal{D}$.
We note that, under the slightly stronger assumption, the discussion in this section can directly be extended to the case where the sets $D_{1},\dots,D_{k}$ are taken from different domains. Specifically, let $\mathcal{D}_{1},\dots,\mathcal{D}_{k}\subseteq 2^{U}$ and assume $(k-1)$-max-distance sparsifiers $\mathcal{K}_{1},\dots,\mathcal{K}_{k}$ of these domains with respect to $2^{U}$ are computed. Then, we can determine whether there exists a $k$-tuple $(D_{1},\dots,D_{k})\in\mathcal{D}_{1}\times\dots\times\mathcal{D}_{k}$ such that $\min_{1\leq i<j\leq k}|D_{i}\triangle D_{j}|\geq d$ (or $\sum_{1\leq i<j\leq k}|D_{i}\triangle D_{j}|\geq d$) by exhaustive search on $\mathcal{K}_{1}\times\dots\times\mathcal{K}_{k}$.
2.2 Clustering
Let $U$ be a finite set, $d\in\mathbb{Z}_{\geq 0}$, $k\in\mathbb{Z}_{\geq 1}$, and $\mathcal{D}\subseteq 2^{U}$. Here, we provide an FPT algorithm for the $k$-center and $k$-sum-of-radii clustering problems using a $(d+1)$-limited $k$-max-distance sparsifier $\mathcal{K}$ of $\mathcal{D}$. For $Z\subseteq U$ and $r\in\mathbb{Z}_{\geq 0}$, the ball of radius $r$ centered at $Z$ is defined as $\mathcal{B}(Z,r):=\{Z^{\prime}\subseteq U\colon|Z\triangle Z^{\prime}|\leq r\}$.
The algorithm first guesses a partition of $\mathcal{K}$ into $k$ clusters $\mathcal{K}_{1},\dots,\mathcal{K}_{k}$. Since $|\mathcal{K}|$ is constant, the cost of this guess is constant. Then, for each $i\in\{1,\dots,k\}$, the algorithm computes the minimum radius $r_{i}$ such that there is a set $D_{i}\in\mathcal{D}$ satisfying $\mathcal{K}_{i}\subseteq\mathcal{B}(D_{i},r_{i})$. If $r_{i}>d$, the algorithm asserts it instead of computing the specific value of $r_{i}$. The $k$-center clustering problem and $k$-sum-of-radii clustering problem on $\mathcal{D}$ are solved by checking whether the maximum and sum, respectively, of the $r_{i}$s is at most $d$. We show the correctness of this algorithm by proving the following.
Lemma 2.4.
Let $(D_{1},\dots,D_{k})\in\mathcal{D}^{k}$ and $(r_{1},\dots,r_{k})\in\{0,\dots,d\}^{k}$. Assume $\mathcal{K}_{i}\subseteq\mathcal{B}(D_{i},r_{i})$ holds for all $i\in\{1,\dots,k\}$. Then, for all $D\in\mathcal{D}$, there is an index $i\in\{1,\dots,k\}$ such that $D\in\mathcal{B}(D_{i},r_{i})$.
Proof.
Assume the contrary. Then, there is a set $D\in\mathcal{D}$ such that for all $i\in\{1,\dots,k\}$, $|D_{i}\triangle D|\geq r_{i}+1$. Since $\mathcal{K}$ is a $(d+1)$-limited $k$-max-distance sparsifier of $\mathcal{D}$, there is a set $K\in\mathcal{K}$ such that for all $i\in\{1,\dots,k\}$, $|D_{i}\triangle K|\geq r_{i}+1$. Hence, $K\not\in\mathcal{K}_{i}$ for all $i\in\{1,\dots,k\}$, contradicting the fact that $(\mathcal{K}_{1},\dots,\mathcal{K}_{k})$ is a partition of $\mathcal{K}$. ∎
We now provide an algorithm to decide whether there exists $D\in\mathcal{D}$ with $\mathcal{K}_{i}\subseteq\mathcal{B}(D,r_{i})$ for each $i\in\{1,\dots,k\}$ and $r_{i}\in\{0,\dots,d\}$. If the domain $\mathcal{D}$ is $2^{U}$, this problem is equivalent to the closest string problem on binary strings, for which a textbook FPT algorithm parameterized by $d$ and $|\mathcal{K}_{i}|$ is known [cygan2015parameterized]. Our algorithm is a modified version of this. An element $e\in U$ is bad if there exist both $K\in\mathcal{K}_{i}$ with $e\in K$ and $K\in\mathcal{K}_{i}$ with $e\not\in K$. The following lemma is fundamental.
Lemma 2.5 ([cygan2015parameterized]).
If there are more than $d|\mathcal{K}_{i}|$ bad elements, no $D\in\mathcal{D}$ satisfies $\mathcal{K}_{i}\subseteq\mathcal{B}(D,d)$.
Let $B$ be the set of bad elements, and assume $|B|\leq d|\mathcal{K}_{i}|$. The algorithm first guesses $B^{\prime}\subseteq B$. The cost of this guess is $2^{d|\mathcal{K}_{i}|}$. Then, it determines whether there exists $D\in\mathcal{D}$ such that $D\cap B=B^{\prime}$ and $\mathcal{K}_{i}\subseteq\mathcal{B}(D,r_{i})$. Let $K^{*}=\mathrm{argmax}_{K\in\mathcal{K}_{i}}|(K\cap B)\triangle B^{\prime}|$. Then, we can claim the following.
Lemma 2.6.
For $D\in\mathcal{D}$ such that $D\cap B=B^{\prime}$, $\max_{K\in\mathcal{K}_{i}}|K\triangle D|=|K^{*}\triangle D|$.
Proof.
Let $K\in\mathcal{K}_{i}$. Then, $|K\triangle D|=|(K\cap B)\triangle(D\cap B)|+|(K\setminus B)\triangle(D\setminus B)|$. From the definition of $B$, the value of $|(K\setminus B)\triangle(D\setminus B)|$ is equal among all $K\in\mathcal{K}_{i}$. Thus, the maximum value of $|K\triangle D|$ for $K\in\mathcal{K}_{i}$ is achieved by the set $K$ that maximizes $|(K\cap B)\triangle(D\cap B)|=|(K\cap B)\triangle B^{\prime}|$. ∎
Now, it is sufficient to solve the problem of determining whether there exists $D\in\mathcal{D}$ such that $D\cap B=B^{\prime}$ and $|K^{*}\triangle D|\leq r_{i}$. This corresponds to the exact extension oracle on $\mathcal{D}$ with $r=r_{i}$, $X=B^{\prime}$, $Y=B\setminus B^{\prime}$, and $C=K^{*}$. Therefore, we can claim the following:
Lemma 2.7.
Assume there exists an FPT algorithm parameterized by $d,k$ that computes a $(d+1)$-limited $k$-max-distance sparsifier of $\mathcal{D}$ with size bounded by a constant that depends only on $d,k$, and the exact extension oracle on $\mathcal{D}$ whose time complexity is FPT parameterized by $r,|X|,|Y|$. Then, there exists an FPT algorithm parameterized by $d,k$ for the $k$-center/$k$-sum-of-radii clustering problem on $\mathcal{D}$.
2.3 Modified Hamming Distance
In some cases, such as when $\mathcal{D}$ is a edge bipartization domain, sets $Z\in\mathcal{D}$ and $U\setminus Z\in\mathcal{D}$ should be considered equivalent. In such cases, it is natural to define the distance between sets $D_{1},D_{2}\in\mathcal{D}$ using the modified Hamming distance $|D_{1}\bar{\triangle}D_{2}|:=\min(|D_{1}\triangle D_{2}|,|D_{1}\triangle(U\setminus D_{2})|)$ instead of the Hamming distance $|D_{1}\triangle D_{2}|$. We can still design FPT algorithms for diversification and clustering problems for this modified Hamming distance using a $d$-limited $k$-max-distance sparsifier. We show that results similar to Lemmas 2.3 and 2.7 also hold for this distance. We assume $(U\setminus D)\in\mathcal{D}$ holds for all $D\in\mathcal{D}$. We prove the following lemmas.
Lemma 2.8.
Assume there exists an FPT algorithm parameterized by $d,k$ that computes a $d$-limited $(2k-2)$-max-distance sparsifier of $\mathcal{D}$ with size bounded by a constant that depends only on $d,k$. Then, there exists an FPT algorithm parameterized by $d,k$ for the max-sum/max-min diversification problem on $\mathcal{D}$ for the modified Hamming distance.
Lemma 2.9.
Assume there exists an FPT algorithm parameterized by $d,k$ that computes a $(d+1)$-limited $2k$-max-distance sparsifier of $\mathcal{D}$ with size bounded by a constant that depends only on $d,k$, and the exact extension oracle on $\mathcal{D}$ whose time complexity is FPT parameterized by $r,|X|,|Y|$. Then, there exists an FPT algorithm parameterized by $d,k$ for the $k$-center/$k$-sum-of-radii clustering problem on $\mathcal{D}$ for the modified Hamming distance.
First, we prove Lemma 2.8. We prove a corresponding result to Lemma 2.1. The remaining part proceeds similarly as in Lemma 2.3.
Lemma 2.10.
Let $U$ be a finite set, $k\in\mathbb{Z}_{\geq 1}$, $d\in\mathbb{Z}_{\geq 0}$, and $\mathcal{D}\subseteq 2^{U}$. Let $\mathcal{K}$ be a $d$-limited $(2k-2)$-max-distance sparsifier of $\mathcal{D}$ and $(D_{1},\dots,D_{k})\in\mathcal{D}^{k}$. Then, there is a $k$-tuple $(K_{1},\dots,K_{k})\in\mathcal{K}^{k}$ such that $\min(d,|D_{i}\bar{\triangle}D_{j}|)\leq\min(d,|K_{i}\bar{\triangle}K_{j}|)$ holds for all $1\leq i<j\leq k$.
Proof.
Assume $(D_{1},\dots,D_{k})\in\mathcal{D}^{k}\setminus\mathcal{K}^{k}$ and let $i$ be an index such that $D_{i}\not\in\mathcal{K}$. It is sufficient to prove that there is a set $K_{i}\in\mathcal{K}$ such that $\min(d,|D_{i}\bar{\triangle}D_{j}|)\leq\min(d,|K_{i}\bar{\triangle}D_{j}|)$ holds for all $j\in\{1,\dots,k\}\setminus\{i\}$. If $(U\setminus D_{i})\in\mathcal{K}$, $K_{i}:=U\setminus D_{i}$ satisfies this condition. Assume otherwise. For $j\in\{1,\dots,k\}\setminus\{i\}$, let $z_{j}:=\min(d,|D_{i}\triangle D_{j}|)$ and $z^{\prime}_{j}:=\min(d,|D_{i}\triangle(U\setminus D_{j})|)$. Since $\mathcal{K}$ is a $d$-limited $(2k-2)$-max-distance sparsifier of $\mathcal{D}$, there exists $K_{i}\in\mathcal{K}$ such that for all $j\in\{1,\dots,k\}\setminus\{i\}$,
$\min(d,|K_{i}\bar{\triangle}D_{j}|)$ $=\min\left(\min(d,|K_{i}\triangle D_{j}|),\min(d,|K_{i}\triangle(U\setminus D_{j})|)\right)\geq\min(\min(d,z_{j}),\min(d,z^{\prime}_{j}))$
$=\min(d,|D_{i}\triangle D_{j}|,|D_{i}\triangle(U\setminus D_{j})|)=\min(d,|D_{i}\bar{\triangle}D_{j}|).\qed$
Next, we prove Lemma 2.9. Let $\mathcal{K}$ be a $(d+1)$-limited $2k$-max-distance sparsifier of $\mathcal{D}$. Similar to Lemma 2.7, the algorithm first guesses a partition of $\mathcal{K}$ into $k$ clusters $\mathcal{K}_{1},\dots,\mathcal{K}_{k}$. Subsequently, for each $i\in\{1,\dots,k\}$, the algorithm determines whether the minimum radius $r_{i}$ of a ball (in the sense of modified Hamming distance) centered at a set of $\mathcal{D}$ that covers $\mathcal{K}_{i}$ is at most $d$, and if so, computes this value. We prove a result corresponding to Lemma 2.4.
Lemma 2.11.
Let $(D_{1},\dots,D_{k})\in\mathcal{D}^{k}$ and $(r_{1},\dots,r_{k})\in\{0,\dots,d\}^{k}$. Assume $|D_{i}\bar{\triangle}K|\leq r_{i}$ holds for all $i\in\{1,\dots,k\}$ and $K\in\mathcal{K}_{i}$. Then, for all $D\in\mathcal{D}$, there is an index $i\in\{1,\dots,k\}$ such that $|D_{i}\bar{\triangle}D|\leq r_{i}$.
Proof.
Assume the contrary. Then, there is a set $D\in\mathcal{D}$ such that for all $i\in\{1,\dots,k\}$, $|D\bar{\triangle}D_{i}|\geq r_{i}+1$, which means $|D\triangle D_{i}|\geq r_{i}+1$ and $|D\triangle(U\setminus D_{i})|\geq r_{i}+1$. Since $\mathcal{K}$ is a $(d+1)$-limited $2k$-max-distance sparsifier of $\mathcal{D}$, there is a set $K\in\mathcal{K}$ such that for all $i\in\{1,\dots,k\}$, $|K\bar{\triangle}D_{i}|=\min(|K\triangle D_{i}|,|K\triangle(U\setminus D_{i})|)\geq r_{i}+1$. Hence, $K\not\in\mathcal{K}_{i}$ for all $i\in\{1,\dots,k\}$, contradicting the fact that $(\mathcal{K}_{1},\dots,\mathcal{K}_{k})$ is a partition of $\mathcal{K}$. ∎
We now provide an algorithm to evaluate the minimum radius of the ball containing $\mathcal{K}_{i}$ in terms of the modified Hamming distance. To determining whether there is $D_{i}\in\mathcal{D}$ such that $|D_{i}\bar{\triangle}K|\leq r_{i}$ for all $K\in\mathcal{K}_{i}$, we guess $\bar{K}\in\{K,U\setminus K\}$ for each $K\in\mathcal{K}_{i}$ and determine whether there is $D_{i}\in\mathcal{D}$ such that $|D_{i}\triangle\bar{K}|\leq r_{i}$ for all $K\in\mathcal{K}_{i}$. After this guessing, the rest are the same as the case discussed in Section 2.2.
3 Framework for $k$-Max-Distance Sparification
In this section, we complete the proof of Theorem 1.1 by providing an FPT algorithm that uses the exact empty extension oracle on $\mathcal{D}$ to obtain a $k$-max-distance sparsifier of $\mathcal{D}$. For further use, we show a slightly more extended result. Let $r\in\mathbb{Z}_{\geq 0}$. We construct a $k$-max-distance sparsifier of $\mathcal{D}$ with respect to $\mathcal{B}(\emptyset,r)$ for $r\geq l$. Since $\mathcal{D}\subseteq\mathcal{B}(\emptyset,l)\subseteq\mathcal{B}(\emptyset,r)$ for $r\geq l$, this is also a $k$-max-distance sparsifier of $\mathcal{D}$ (with respect to $\mathcal{D}$).
A set family $\mathcal{S}:=\{S_{1},\dots,S_{t}\}$ is called a sunflower if there exists a set called core $C$ such that for any $1\leq i<j\leq t$, $S_{i}\cap S_{j}=C$. The following is well-known.
Lemma 3.1 (Sunflower Lemma [cygan2015parameterized, erdos1960intersection]).
Let $U$ be a finite set, $l,t\in\mathbb{Z}_{\geq 0}$, and $\mathcal{K}\subseteq 2^{U}$ be a family consisting only of sets of size at most $l$. If $|\mathcal{K}|>l!(t-1)^{l}$, then $\mathcal{K}$ contains a sunflower of size $t$.
For $t\in\mathbb{Z}_{\geq 0}$, $\mathcal{T}\subseteq 2^{U}$, and $Z\in 2^{U}\setminus\mathcal{T}$, a sunflower $\mathcal{S}\subseteq\mathcal{T}$ is a $(Z,t)$-sunflower of $\mathcal{T}$ if it satisfies the following three conditions.
  • $|\mathcal{S}|=t$,
  • For each $S\in\mathcal{S}$, $|S|=|Z|$, and
  • The core of $\mathcal{S}$ is a subset of $Z$.
The following lemma is the core of our framework.
Lemma 3.2.
Let $U$ be a finite set, $\mathcal{D}\subseteq 2^{U}$, and $\mathcal{K}\subseteq\mathcal{D}$ be a $k$-max-distance sparsifier of $\mathcal{D}$ with respect to $\mathcal{B}(\emptyset,r)$. Let $Z\in\mathcal{K}$ and assume there is a $(Z,kr+1)$-sunflower $\mathcal{S}$ of $\mathcal{K}\setminus\{Z\}$. Then, $\mathcal{K}\setminus\{Z\}$ is also a $k$-max-distance sparsifier of $\mathcal{D}$ with respect to $\mathcal{B}(\emptyset,r)$.
Proof.
Let $(F_{1},\dots,F_{k})\in\mathcal{B}(\emptyset,r)^{k}$ and $(z_{1},\dots,z_{k})\in\mathbb{Z}_{\geq 0}^{k}$. We show the equivalence of the following two conditions:
  • There exists $K\in\mathcal{K}$ such that for each $i\in\{1,\dots,k\}$, $|F_{i}\triangle K|\geq z_{i}$.
  • There exists $K\in\mathcal{K}\setminus\{Z\}$ such that for each $i\in\{1,\dots,k\}$, $|F_{i}\triangle K|\geq z_{i}$.
Since $\mathcal{K}\setminus\{Z\}\subseteq\mathcal{K}$, the latter implies the former. We assume the former and prove the latter. Take the $K$ that satisfies the former condition. If $K\neq Z$, the claim is obvious, so assume $K=Z$. Then, there exists a $(Z,kr+1)$-sunflower $\mathcal{S}$ of $\mathcal{K}\setminus\{Z\}$. Let $C$ be the core of $\mathcal{S}$. Since $\left|\bigcup_{i\in\{1,\dots,k\}}F_{i}\right|\leq\sum_{i\in\{1,\dots,k\}}|F_{i}|\leq kr$, there exists an $S\in\mathcal{S}$ such that $F_{i}\cap(S\setminus C)=\emptyset$ for all $i\in\{1,\dots,k\}$. In this case, for each $i\in\{1,\dots,k\}$, $|F_{i}\triangle S|=|F_{i}|+|S|-2|F_{i}\cap S|=|F_{i}|+|Z|-2|F_{i}\cap C|\geq|F_{i}|+|Z|-2|F_{i}\cap Z|=|F_{i}\triangle Z|\geq z_{i}$, where the first inequality follows from $C\subseteq Z$. Thus, the lemma is proved. ∎
1 Procedure KSparsify($k,r$)
       Input: $k\in\mathbb{Z}_{\geq 1}$, $r\in\mathbb{Z}_{\geq 0}$
2       Let $\mathcal{K}:=\emptyset$;
3       Let $f:=\mathsf{true}$;
4       while $f=\mathsf{true}$ do
5             Let $R:=\bigcup_{K\in\mathcal{K}}K$;
6             Let $f^{\prime}:=\mathsf{false}$;
7             for $l^{\prime}\in\{0,\dots,l\}$ do
8                   Let $\mathfrak{S}$ be the family of all sunflowers $\mathcal{S}\subseteq\mathcal{K}$ such that $|\mathcal{S}|=kr+1$ and each $S\in\mathcal{S}$ satisfies $|S|=l^{\prime}$;
9                   for $Y\subseteq R$ that intersects with all $K\in\mathcal{K}$ with $|K|=l^{\prime}$ and the cores of all sunflowers of $\mathfrak{S}$ do
10                         Let $D=\textsc{ExactEmptyExtension}(l^{\prime},Y)$;
11                         if $D\neq\bot$ and $f^{\prime}=\mathsf{false}$ then
12                               Add $D$ to $\mathcal{K}$;
13                               $f^{\prime}:=\mathsf{true}$;
14
15
16
17            if $f^{\prime}=\mathsf{false}$ then
18                   $f:=\mathsf{false}$;
19
20
21      return $\mathcal{K}$;
22
Algorithm 1 $k$-max-distance sparsification of $\mathcal{D}$
Our algorithm is given in Algorithm 1, where ExactEmptyExtension($l^{\prime},Y$) represents the exact empty extension oracle on $\mathcal{D}$ with arguments $l^{\prime}$ and $Y$. The algorithm starts with $\mathcal{K}:=\emptyset$ and repeatedly adds $Z\in\mathcal{D}\setminus\mathcal{K}$ such that there is no $(Z,kr+1)$-sunflower of $\mathcal{K}$ to $\mathcal{K}$. The following lemma shows this algorithm stops after a constant number of iterations.
Lemma 3.3.
The number of iterations of the loop starting from line 1 in Algorithm 1, as well as the size of the output family, is at most $(l+1)!(kr+1)^{l}$.
Proof.
By the definition of $Y$ in line 1, the $D$ added to $\mathcal{K}$ in line 1 is distinct from any set in $\mathcal{K}$, since $Y$ has a nonempty intersection with each $D\in\mathcal{K}$ with $|D|=l^{\prime}$. Moreover, there is no $(D,kr+1)$-sunflower in $\mathcal{K}$, since $Y$ has a nonempty intersection with the core of any sunflower in $\mathfrak{S}$. Thus, adding $D$ to $\mathcal{K}$ does not form a new sunflower of size $kr+2$ consisting of sets of size $|D|$. Hence, by Lemma 3.1, for each $l^{\prime}\in\{0,\dots,l\}$, the number of sets of size $l^{\prime}$ in $\mathcal{K}$ is at most $l!(kr+1)^{l}$. Thus, the number of iterations of the loop and the size of the output family is at most $(l+1)!(kr+1)^{l}$. ∎
In particular, at each step of the algorithm, since $|R|\leq|\mathcal{K}|l\leq(l+1)!(kr+1)^{l}l$, the size of $Y$ chosen in line 1 is bounded by a constant. The time complexity is bounded as follows.
Lemma 3.4.
Algorithm 1 makes at most $2^{2^{O(l\log(klr))}}$ calls of ExactEmptyExtension($\cdot$) and has a time complexity of $2^{2^{O(l\log(klr))}}$.
Proof.
By Lemma 3.3, the number of iterations of the loop starting from line 1 is $2^{O(l\log(klr))}$. Therefore, at each step of the algorithm, $|R|\leq 2^{O(l\log(klr))}$. We can compute $\mathfrak{S}$ by exhaustively checking subsets of $\mathcal{K}$ of size $kr+1$, which takes $|\mathcal{K}|^{kr+1}\leq 2^{O(klr\log(klr))}$ time. The number of candidates for $Y$ in line 1 is at most $2^{|R|}\leq 2^{2^{O(l\log klr)}}$, which dominates the time complexity. ∎
The correctness of the algorithm is shown as follows.
Lemma 3.5.
Algorithm 1 outputs a $k$-max-distance sparsifier of $\mathcal{D}$ with respect to $\mathcal{B}(\emptyset,r)$.
Proof.
Let $\mathcal{K}$ be the output of Algorithm 1 and $\mathcal{D}\setminus\mathcal{K}=\{D_{1},\dots,D_{|\mathcal{D}\setminus\mathcal{K}|}\}$. By the termination condition of the algorithm, for each $i\in\{1,\dots,|\mathcal{D}\setminus\mathcal{K}|\}$, there exists a $(D_{i},kr+1)$-sunflower of $\mathcal{K}$. By Lemma 3.2, for each $i\in\{1,\dots,|\mathcal{D}\setminus\mathcal{K}|\}$, if $\mathcal{D}\setminus\{D_{1},\dots,D_{i-1}\}$ is a $k$-max-distance sparsifier of $\mathcal{D}$ with respect to $\mathcal{B}(\emptyset,r)$, then $\mathcal{D}\setminus\{D_{1},\dots,D_{i}\}$ also is. Therefore, $\mathcal{K}=\mathcal{D}\setminus\{D_{1},\dots,D_{|\mathcal{D}\setminus\mathcal{K}|}\}$ is a $k$-max-distance sparsifier of $\mathcal{D}$ with respect to $\mathcal{B}(\emptyset,r)$. ∎
4 Framework for $d$-Limited $k$-Max-Distance Sparsification
4.1 Overall Flow
In this section, we complete the proof of Theorem 1.2 by providing FPT algorithm that uses the $(-1,1)$-optimization oracle and the exact extension oracle on $\mathcal{D}$ to obtain a $d$-limited $k$-max-distance sparsifier of $\mathcal{D}$. Actually, for further applications, we construct the slightly more general object of $d$-limited $k$-max-distance sparsifier of $\mathcal{D}$ with respect to $2^{U}$, not with respect to $\mathcal{D}$ itself. Our framework is given by Algorithm 2, which consists of two steps. Let $p\in\mathbb{Z}_{\geq 0}$ be an integer with $2d<p$. The first step achieves one of the following.
  • Find a set $\mathcal{C}\subseteq\mathcal{D}$ with size at most $k$ such that $\mathcal{D}\subseteq\bigcup_{C\in\mathcal{C}}\mathcal{B}(C,p)$.
  • Find a set $\mathcal{C}\subseteq\mathcal{D}$ of size $k+1$ such that $|C\triangle C^{\prime}|>2d$ holds for any distinct $C,C^{\prime}\in\mathcal{C}$.
We do this by using the following approximate far set oracle, which will be designed in Section 4.2.
1 Procedure DKSparsify($k,d,p$)
       Input: $k\in\mathbb{Z}_{\geq 1}$, $d,p\in\mathbb{Z}_{\geq 0}$ such that $2d<p$
2       Let $\mathcal{C}\leftarrow\emptyset$;
3       while true do
4             Find a set $D\in\mathcal{D}$ such that $|D\triangle C|\geq 2d$ for all $C\in\mathcal{C}$, or correctly conclude that $\bigcup_{C\in\mathcal{C}}(\mathcal{D}\cap\mathcal{B}(C,p))=\mathcal{D}$;
5             if Such $D$ is found then
6                   Add $D$ to $\mathcal{C}$;
7            else
8                   break;
9            if $|\mathcal{C}|>k$ then
10                   return $\mathcal{C}$;
11
12
13      Let $\mathcal{K}\leftarrow\emptyset$;
14       for $C\in\mathcal{C}$ do
15             Compute $k$-max-distance sparsifier $\mathcal{K}^{*}_{C}$ of $\mathcal{D}^{*}_{C}:=\{D\triangle C\colon D\in\mathcal{D}_{C}:=\mathcal{D}\cap\mathcal{B}(C,p)\}$ with respect to $\mathcal{B}(C,p+d)$;
16             $\mathcal{K}\leftarrow\mathcal{K}\cup\{K^{*}\triangle C\colon K^{*}\in\mathcal{K}^{*}_{C}\}$;
17
18      return $\mathcal{K}$;
19
Algorithm 2 The framework for $d$-limited $k$-max-distance sparsifiction
Approximate Far Set Oracle: Let $U$ be a finite set, $d\in\mathbb{Z}_{\geq 0}$, $\mathcal{D}\subseteq 2^{U}$, and $\mathcal{C}\subseteq\mathcal{D}$. The approximate far set oracle returns one of the following.
  • A set of $\mathcal{D}$ that does not belong to $\bigcup_{C\in\mathcal{C}}\mathcal{B}(C,2d)$.
  • $\bot$. This option can be chosen only when $\mathcal{D}\subseteq\bigcup_{C\in\mathcal{C}}\mathcal{B}(C,p)$.
Starting with $\mathcal{C}:=\emptyset$, we repeat the following steps. If the approximate far set oracle returns $\bot$, terminate the loop. Otherwise, add the element found by the oracle to $\mathcal{C}$. If the oracle returns $\bot$ within $k$ iterations, the first condition is achieved. If not, the set $\mathcal{C}$ after $k+1$ iterations satisfies the second condition. In the latter case, the following lemma shows that $\mathcal{C}$ is a $d$-limited $k$-max-distance sparsifier.
Lemma 4.1.
Let $r\in\mathbb{Z}_{\geq 0}$. Let $\mathcal{C}$ be a subset of $\mathcal{D}$ of size $k+1$ such that for any distinct $C,C^{\prime}\in\mathcal{C}$, $|C\triangle C^{\prime}|\geq 2d$. Then, $\mathcal{C}$ is a $d$-limited $k$-max-distance sparsifier of $\mathcal{D}$ with respect to $2^{U}$.
Proof.
For each $F\subseteq U$, there is at most one $C\in\mathcal{C}$ such that $|F\triangle C|<d$. Hence, for any $k$-tuple $(F_{1},\dots,F_{k})$ of the subsets of $U$, there exists a $C\in\mathcal{C}$ such that $|F_{i}\triangle C|\geq d$ for all $i\in\{1,\dots,k\}$. In particular, for all $(z_{1},\dots,z_{k})\in\{0,\dots,d\}^{k}$, the following two conditions are always true and thus equivalent.
  • There exists $D\in\mathcal{D}$ such that for each $i\in\{1,\dots,k\}$, $|F_{i}\triangle D|\geq z_{i}$.
  • There exists $C\in\mathcal{C}$ such that for each $i\in\{1,\dots,k\}$, $|F_{i}\triangle C|\geq z_{i}$.∎
Now, we assume the first condition. Let $\mathcal{C}$ be a subset of $\mathcal{D}$ of size at most $k$ such that $\mathcal{D}\subseteq\bigcup_{C\in\mathcal{C}}\mathcal{B}(C,p)$. The second step involves constructing a $d$-limited $k$-max-distance sparsifier of $\mathcal{D}_{C}:=\mathcal{D}\cap\mathcal{B}(C,p)$ with respect to $\mathcal{B}(C,p+d)$ for each $C\in\mathcal{C}$. We prove that the union of all such $d$-limited $k$-max-distance sparsifiers obtained in this manner is a $d$-limited $k$-max-distance sparsifier of $\mathcal{D}$ with respect to $2^{U}$.
Lemma 4.2.
Assume $\mathcal{D}=\bigcup_{C\in\mathcal{C}}\mathcal{D}_{C}$. For each $C\in\mathcal{C}$, let $\mathcal{K}_{C}\subseteq\mathcal{D}_{C}$ be a $d$-limited $k$-max-distance sparsifier of $\mathcal{D}_{C}$ with respect to $\mathcal{B}(C,p+d)$. Then, $\mathcal{K}:=\bigcup_{C\in\mathcal{C}}\mathcal{K}_{C}$ is a $d$-limited $k$-max-distance sparsifier of $\mathcal{D}$ with respect to $2^{U}$.
Proof.
Let $(F_{1},\dots,F_{k})\in\left(2^{U}\right)^{k}$ and $(z_{1},\dots,z_{k})\in\{0,\dots,d\}^{k}$. We show the equivalence of the following two conditions.
  • There exists $D\in\mathcal{D}$ such that for each $i\in\{1,\dots,k\}$, $|F_{i}\triangle D|\geq z_{i}$.
  • There exists $K\in\mathcal{K}$ such that for each $i\in\{1,\dots,k\}$, $|F_{i}\triangle K|\geq z_{i}$.
Since $\mathcal{K}\subseteq\mathcal{D}$, the latter implies the former. We assume the former and prove the latter. Take any $D\in\mathcal{D}$ that satisfies the former condition. Since $\mathcal{D}=\bigcup_{C\in\mathcal{C}}\mathcal{D}_{C}$, we can take $C\in\mathcal{C}$ such that $D\in\mathcal{D}_{C}$. Now, define $(F^{\prime}_{1},\dots,F^{\prime}_{k})\in\mathcal{B}(C,p+d)^{k}$ and $(z^{\prime}_{1},\dots,z^{\prime}_{k})\in\{0,\dots,p\}^{k}$ as follows.
$F^{\prime}_{i}=\begin{cases}F_{i}&\text{if }F_{i}\in\mathcal{B}(C,p+d)\\ C&\text{otherwise}\end{cases},\quad\quad z^{\prime}_{i}=\begin{cases}z_{i}&\text{if }F_{i}\in\mathcal{B}(C,p+d)\\ 0&\text{otherwise}\end{cases}$
Clearly, for each $i\in\{1,\dots,k\}$, $F^{\prime}_{i}\in\mathcal{B}(C,p+d)$. Furthermore, if $F_{i}\in\mathcal{B}(C,p+d)$, then $|F^{\prime}_{i}\triangle D|=|F_{i}\triangle D|\geq z_{i}=z^{\prime}_{i}$. If $F_{i}\not\in\mathcal{B}(C,p+d)$, $|F^{\prime}_{i}\triangle D|\geq 0=z^{\prime}_{i}$. Therefore, for each $i\in\{1,\dots,k\}$, $|F^{\prime}_{i}\triangle D|\geq z^{\prime}_{i}$. Since $\mathcal{K}_{C}$ is a $d$-limited $k$-max-distance sparsifier of $\mathcal{D}_{C}$ with respect to $\mathcal{B}(C,p+d)$, there exists $K\in\mathcal{K}_{C}\subseteq\mathcal{K}$ such that for each $i\in\{1,\dots,k\}$, $|F^{\prime}_{i}\triangle K|\geq z^{\prime}_{i}$. We show that this $K$ satisfies the latter condition. For each $i\in\{1,\dots,k\}$, we show that $|F_{i}\triangle K|\geq z_{i}$. If $F_{i}\in\mathcal{B}(C,p+d)$, then $|F_{i}\triangle K|=|F^{\prime}_{i}\triangle K|\geq z^{\prime}_{i}=z_{i}$. If $F_{i}\not\in\mathcal{B}(C,p+d)$, since $K\in\mathcal{K}_{C}\subseteq\mathcal{D}_{C}\subseteq\mathcal{B}(C,p)$, we have $|F_{i}\triangle K|\geq|F_{i}\triangle C|-|K\triangle C|>(p+d)-p\geq d\geq z_{i}$, where the first inequality is from the triangle inequality. ∎
Next, we reduce the computation of $d$-limited $k$-max-distance sparsifiers to the computation of $k$-max-distance sparsifiers of families consisting of constant-size sets, which was discussed in Section 3. For $C\in\mathcal{C}$, let $\mathcal{D}_{C}^{*}:=\{D\triangle C\mid D\in\mathcal{D}_{C}\}$. By the definition of $\mathcal{D}_{C}$, we have $\mathcal{D}_{C}^{*}\subseteq\mathcal{B}(\emptyset,p)$. The following holds.
Lemma 4.3.
Let $C\in\mathcal{C}$. A subset $\mathcal{K}_{C}\subseteq\mathcal{D}_{C}$ is a $d$-limited $k$-max-distance sparsifier of $\mathcal{D}_{C}$ with respect to $\mathcal{B}(C,p+d)$ if and only if $\mathcal{K}_{C}^{*}:=\{K\triangle C\mid K\in\mathcal{K}_{C}\}$ is a $d$-limited $k$-max-distance sparsifier of $\mathcal{D}_{C}^{*}$ with respect to $\mathcal{B}(\emptyset,p+d)$.
Proof.
Let $(F_{1},\dots,F_{k})\in\mathcal{B}(C,p+d)^{k}$ and $(z_{1},\dots,z_{k})\in\{0,\dots,d\}^{k}$. For each $i\in\{1,\dots,k\}$, let $F_{i}^{*}:=F_{i}\triangle C$. It follows that $(F_{1}^{*},\dots,F_{k}^{*})\in\mathcal{B}(\emptyset,p+d)^{k}$. Since $|F_{i}\triangle D|=|(F_{i}\triangle C)\triangle(D\triangle C)|=|F_{i}^{*}\triangle(D\triangle C)|$, the following two conditions are equivalent.
  • (i)
    There exists $D\in\mathcal{D}_{C}$ such that for each $i\in\{1,\dots,k\}$, $|F_{i}\triangle D|\geq z_{i}$.
  • (i’)
    There exists $D^{*}\in\mathcal{D}_{C}^{*}$ such that for each $i\in\{1,\dots,k\}$, $|F_{i}^{*}\triangle D^{*}|\geq z_{i}$.
Similarly, the following two conditions are also equivalent.
  • (ii)
    There exists $K\in\mathcal{K}_{C}$ such that for each $i\in\{1,\dots,k\}$, $|F_{i}\triangle K|\geq z_{i}$.
  • (ii’)
    There exists $K^{*}\in\mathcal{K}_{C}^{*}$ such that for each $i\in\{1,\dots,k\}$, $|F_{i}^{*}\triangle K^{*}|\geq z_{i}$.
Hence, (i) and (ii) are equivalent if and only if (i’) and (ii’) are equivalent, and the lemma is proved. ∎
If $\mathcal{K}_{C}^{*}$ is a $k$-max-distance sparsifier of $\mathcal{D}_{C}^{*}$ with respect to $\mathcal{B}(\emptyset,p+d)$, then it is also a $d$-limited $k$-max-distance sparsifier of $\mathcal{D}_{C}^{*}$ with respect to $\mathcal{B}(\emptyset,p+d)$ for any $d\in\mathbb{Z}_{\geq 0}$. Therefore, a $d$-limited $k$-max-distance sparsifier $\mathcal{K}$ of $\mathcal{D}$ with respect to $\mathcal{B}(C,p+d)$ can be computed as
$\mathcal{K}:=\bigcup_{C\in\mathcal{C}}\{K^{*}\triangle C\colon K^{*}\in\mathcal{K}^{*}_{C}\}.$
From the discussion in Section 3, $\mathcal{K}_{C}^{*}$ can be obtained by calling the exact empty extension oracle on $\mathcal{D}_{C}^{*}$ a constant number of times that depends on $k$, $l=p$, and $r=p+d$. The exact empty extension oracle for $\mathcal{D}_{C}^{*}$ is equivalent to the exact extension oracle on $\mathcal{D}_{C}$ when the inputs $C,X,Y$ are taken to be $C,Y\cap C,Y\setminus C$, respectively. Therefore, $\mathcal{K}_{C}^{*}$ can be obtained by calling the exact extension oracle on $\mathcal{D}_{C}$ a constant number of times that depends only on $k$ and $p$.
4.2 Designing the Approximate Far Set Oracle
Here, we design a randomized algorithm parameterized by $|\mathcal{C}|$ and $d$ for the approximate far set oracle. The algorithm is given in Algorithm 3. Our algorithm repeats the following sufficient number of times: It selects a weight vector $w\in\{-1,1\}^{U}$ uniformly at random and finds a set $D\in\mathcal{D}$ that maximizes $w(D):=\sum_{e\in D}w_{e}$. If the found $D$ does not belong to $\bigcup_{C\in\mathcal{C}}\mathcal{B}(C,2d)$, it outputs this $D$ and terminates. If no such $D$ is found after a sufficient number of iterations, it returns $\bot$. We now prove the correctness of the algorithm. We can claim the following.
Lemma 4.4.
Assume $\max_{D\in\mathcal{D}}w(D)>\max_{C\in\mathcal{C}}w(C)+2d$. Then, the $D$ that attains the maximum on the left-hand side does not belong to $\bigcup_{C\in\mathcal{C}}\mathcal{B}(C,2d)$.
Proof.
For each $C\in\mathcal{C}$, $|D\triangle C|=\sum_{e\in D\triangle C}|w_{e}|\geq w(D)-w(C)>2d$. ∎
1 Procedure ApproxFarSet($\mathcal{C},k,d$)
       Input: $\mathcal{C}\subseteq 2^{U}$, $k\in\mathbb{Z}_{\geq 1}$, $d\in\mathbb{Z}_{\geq 0}$
2       for $i\in\left\{1,\dots,2^{2^{O(k)}}\right\}$ do
3             Sample weight $w\in\{-1,1\}^{U}$ uniformly at random;
4             Let $D:=\mathrm{argmax}_{D\in\mathcal{D}}w(D)$;
5             if $|D\triangle C|>2d$ for all $C\in\mathcal{C}$ then
6                   return $D$;
7
8
9      return $\bot$;
10
Algorithm 3 Algorithm for approximate far set oracle
The following lemma is the core of the analysis.
Lemma 4.5.
Assume $p\geq(4d+2)^{2}\cdot 2^{k-1}$ and $|\mathcal{C}|\leq k$. Let $D\in\mathcal{D}$ and assume $D\not\in\bigcup_{C\in\mathcal{C}}\mathcal{B}(C,p)$. Then,
$\Pr\left[w(D)>\max_{C\in\mathcal{C}}w(C)+2d\right]\geq 2^{-2^{O(k)}}.$
Proof.
Let $\mathcal{C}^{*}:=\{C\triangle D\colon C\in\mathcal{C}\}$ and $w^{*}\in\{-1,1\}^{U}$ be a weight vector such that $w^{*}_{e}=w_{e}$ for $e\in D$ and $w^{*}_{e}=-w_{e}$ otherwise. Then, for all $C\in\mathcal{C}$, $w(C)=w(D)-w(D\setminus C)+w(C\setminus D)=w(D)-w^{*}(C\triangle D)$. Therefore, it is sufficient to prove that
$\Pr\left[\min_{C^{*}\in\mathcal{C}^{*}}w^{*}(C^{*})>2d\right]\geq 2^{-2^{O(k)}}.$
Since $D\not\in\bigcup_{C\in\mathcal{C}}\mathcal{B}(C,p)$, $|C^{*}|>p$ holds for all $C^{*}\in\mathcal{C}^{*}$. Moreover, $w^{*}$ follows a uniform distribution over $\{-1,1\}^{U}$.
Two elements $e,e^{\prime}\in U$ are equivalent under $\mathcal{C}^{*}$ if $e\in C^{*}$ and $e^{\prime}\in C^{*}$ are equivalent for all $C^{*}\in\mathcal{C}^{*}$. We partition the ground set $U$ into $2^{|\mathcal{C}^{*}|}$ subsets such that elements in the same subset are equivalent under $\mathcal{C}^{*}$. Specifically, for each $\mathcal{X}\subseteq\mathcal{C}^{*}$, define $U_{\mathcal{X}}:=\{e\in U\colon\{C^{*}\in\mathcal{C}^{*}\colon C^{*}\ni e\}=\mathcal{X}\}$. For each $C^{*}\in\mathcal{C}^{*}$, define $\mathcal{X}_{C^{*}}\subseteq\mathcal{C}^{*}$ as the family where $|U_{\mathcal{X}}|$ is maximized among all subsets $\mathcal{X}\subseteq\mathcal{C}^{*}$ with $C^{*}\in\mathcal{X}$. Each $C^{*}\in\mathcal{C}^{*}$ satisfies $|C^{*}|>p$, and there are $2^{|\mathcal{C}^{*}|-1}$ subfamilies $\mathcal{X}\subseteq\mathcal{C}^{*}$ with $C^{*}\in\mathcal{X}$, thus by the pigeonhole principle, $|U_{\mathcal{X}}|>\frac{p}{2^{|\mathcal{C}^{*}|-1}}=\frac{p}{2^{|\mathcal{C}|-1}}\geq(4d+2)^{2}$.
Now, assume the weight vector $w^{*}$ satisfies the following two conditions.
  • For $\mathcal{X}\subseteq\mathcal{C}^{*}$ such that $\mathcal{X}=\mathcal{X}_{C^{*}}$ holds for some $C^{*}\in\mathcal{C}^{*}$, |$\left|\{e\in U_{\mathcal{X}}\colon w^{*}_{e}=1\}\right|>\frac{|U_{\mathcal{X}}|}{2}+d$.
  • For all other $\mathcal{X}\subseteq\mathcal{C}^{*}$, |$\left|\{e\in U_{\mathcal{X}}\colon w^{*}_{e}=1\}\right|\geq\frac{|U_{\mathcal{X}}|}{2}$.
Then, for each $C^{*}\in\mathcal{C}^{*}$,
$w^{*}(C^{*})=\sum_{\mathcal{X}\ni C^{*}}\left(\left|\{e\in U_{\mathcal{X}}\colon w^{*}_{e}=1\}\right|-\left|\{e\in U_{\mathcal{X}}\colon w^{*}_{e}=-1\}\right|\right)>0\cdot(2^{|\mathcal{C}^{*}|-1}-1)+2d\cdot 1=2d.$
Thus, we need to show that the probability that $w^{*}$ satisfies these two conditions is at least $2^{-2^{O(k)}}$. Each of the conditions for $\mathcal{X}\subseteq\mathcal{C}^{*}$ is independent, so we evaluate the probability for a fixed $\mathcal{X}$. The probability that a fixed $\mathcal{X}$ satisfies the second condition is clearly at least $\frac{1}{2}$. Assume there exists $C^{*}\in\mathcal{C}^{*}$ with $\mathcal{X}=\mathcal{X}_{C^{*}}$. Let $N:=|U_{\mathcal{X}}|>(4d+2)^{2}$. Then,
$\Pr\left[\left|\{e\in U_{\mathcal{X}}\colon w^{*}_{e}=1\}\right|>\frac{N}{2}+d\right]$ $=\frac{1}{2}\cdot\left(1-\frac{\sum_{p=\left\lceil{\frac{N}{2}-d}\right\rceil}^{\left\lfloor{\frac{N}{2}+d}\right\rfloor}\binom{N}{p}}{2^{N}}\right)$
$\geq\frac{1}{2}\cdot\left(1-\frac{(2d+1)\cdot\frac{2^{N}}{\sqrt{N}}}{2^{N}}\right)$
$\geq\frac{1}{2}\cdot\left(1-\frac{2d+1}{4d+2}\right)=\frac{1}{4},$
where the first inequality follows from $\binom{N}{p}\leq\frac{2^{N}}{\sqrt{N}}$. Thus, the probability that $w^{*}$ satisfies the two conditions is at least $2^{-2^{|\mathcal{C}^{*}|}}\cdot 4^{-|\mathcal{C}^{*}|}\geq 2^{-2^{O(k)}}$. ∎
By repeating the sampling of $w$ a sufficient number of times, we can state the following.
Lemma 4.6.
Let $\epsilon>0$, $\mathcal{D},\mathcal{C}\subseteq 2^{U}$, $k\in\mathbb{Z}_{\geq 1}$, and $d\in\mathbb{Z}_{\geq 0}$. Assume $|\mathcal{C}|\leq k$. Then, there exists a randomized algorithm that runs in time $2^{2^{O(k)}}\log\epsilon^{-1}$ and satisfies the following.
  • If there exists $D\in\mathcal{D}$ such that $D\not\in\bigcup_{C\in\mathcal{C}}\mathcal{B}(C,(4d+2)^{2}\cdot 2^{k})$, the algorithm returns a set $D^{\prime}\in\mathcal{D}$ satisfying $D^{\prime}\not\in\bigcup_{C\in\mathcal{C}}\mathcal{B}(C,2d)$ with probability at least $1-\epsilon$.
  • If not, the algorithm returns either $\bot$ or a set $D^{\prime}\in\mathcal{D}$ satisfying $D^{\prime}\not\in\bigcup_{C\in\mathcal{C}}\mathcal{B}(C,2d)$.
Combining Lemma 4.6 with the results from Sections 3 and 4.1, we have the following.
Lemma 4.7.
Let $\epsilon>0$, $\mathcal{D}\subseteq 2^{U}$, $k\in\mathbb{Z}_{\geq 1}$, and $d\in\mathbb{Z}_{\geq 0}$. Then, there exists a randomized algorithm that, with probability $1-\epsilon$, computes a $d$-limited $k$-max-distance sparsifier of $\mathcal{D}$ with respect to $2^{U}$ with size at most $2^{2^{O(k+\log d)}}$ in time $\left(2^{2^{2^{O(k+\log d)}}}+2^{2^{O(k)}}\log\epsilon^{-1}\right)poly(|U|)$. It calls at most $2^{2^{O(k)}}\log\epsilon^{-1}$ calls to the $(-1,1)$-optimization oracle on $\mathcal{D}$ and at most $2^{2^{2^{O(k+\log d)}}}$ calls to the exact extension oracle on $\mathcal{D}$ such that $r\leq(4d+2)^{2}2^{k-1}$ and $|X|,|Y|\leq 2^{2^{O(k+\log d)}}$.
Proof.
Since the algorithm calls the approximate far set oracle at most $k+1$ times, applying Lemma 4.6 with an error $\frac{\epsilon}{k+1}$ claims that the algorithm computes $\mathcal{C}$ correctly with probability $1-\epsilon$ by making $2^{2^{O(k)}}\log\epsilon^{-1}$ calls of the $(-1,1)$-optimization oracle. If $\mathcal{C}$ is correctly computed, applying Algorithm 1 on $\mathcal{D}^{*}_{C}$ for each $C\in\mathcal{C}$ computes the $d$-limited $k$-max-distance sparsifier of $\mathcal{D}$ with respect to $2^{U}$. Applying Lemma 3.3 for $l=p$ and $r=p+d$ bounds the size of the output sparsifier by $(p+1)!(k(p+d)+1)^{p}\cdot k\leq 2^{O(p\log(kp(p+d)))}\leq 2^{O(2^{k}d^{2}\log(4^{d}d^{4}k))}\leq 2^{2^{O(k+\log d)}}$. Applying Lemma 3.4 for $l=p$ and $r=p+d$ bounds the number of calls to the exact extension oracle by $2^{2^{O(p\log(kp(p+d)))}}\leq 2^{2^{O(2^{k}d^{2}\log(4^{d}d^{4}k))}}\leq 2^{2^{2^{O(k+\log d)}}}$. Moreover, for each call of the exact extension oracle, we have $r\leq p=(4d+2)^{2}2^{k-1}$ and $|X|,|Y|\leq 2^{2^{O(k+\log d)}}\cdot p\leq 2^{2^{O(k+\log d)}}$. ∎
5 Application
In this section, we show that Theorems 1.1 and 1.2 actually yield algorithms for diversification and clustering problems on several domains by constructing the required oracles for these domains.
We remark on a simple but useful fact. Assume a domain $\mathcal{D}$ can be written as the union of domains $\mathcal{D}_{1},\dots,\mathcal{D}_{t}$. Then, all oracles required by Theorems 1.1 or 1.2 for the domain $\mathcal{D}$ can be implemented by calling a corresponding oracle for each domain $\mathcal{D}_{1},\dots,\mathcal{D}_{t}$. Specifically, the $(-1,1)$-optimization oracle on $\mathcal{D}$ can be realized by returning the maximum of the outputs among each call, and the exact (empty) extension oracle on $\mathcal{D}$ can be realized by returning any solution returned by a call if any call returns a $\bot$ solution. Henceforth, for any domain that can be written as a union of an FPT number of domains to which we provide FPT algorithms using our framework, our framework automatically yields the same results.
5.1 $k,l$ as Parameters
Here we apply Theorem 1.1 to specific domains $\mathcal{D}$. We design an exact empty extension oracle on $\mathcal{D}$ that runs in FPT time. As mentioned in Section 1.2.2, in most cases, the exact empty extension oracle on $\mathcal{D}$ can be designed almost directly from an algorithm that finds a set of size $r$ in $\mathcal{D}$. Therefore, most designs for the exact empty extension oracle in this section are straightforward.
5.1.1 Vertex Cover
Let $G=(V,E)$ be an undirected graph and $l\in\mathbb{Z}_{\geq 0}$. A vertex subset $Z\subseteq V$ is a vertex cover of $G$ if for each $e\in E$, $e\cap Z\neq\emptyset$. Let $\mathcal{D}\subseteq 2^{V}$ be the set of all vertex covers of $G$ of size at most $l$. [baste2019fpt] provided an FPT algorithm parameterized by $k,l$ for the max-min/max-sum diversification problems on $\mathcal{D}$.
We design an exact empty extension oracle on $\mathcal{D}$. Let $r\in\mathbb{Z}_{\geq 0}$ and $Y\subseteq V$. Define $\mathcal{D}_{Y,r}$ as the family of all vertex covers of size $r$ of $G$ disjoint from $Y$. We construct an algorithm to find a set in $\mathcal{D}_{Y,r}$. If $Y$ includes two adjacent vertices, $\mathcal{D}_{Y,r}$ is obviously empty. Otherwise, let $Z$ be the set of vertices adjacent to at least one vertex in $Y$, and let $\mathcal{D}^{\prime}$ be the family of all vertex covers of size $r-|Z|$ of the subgraph $G^{\prime}$ induced by $V\setminus(Y\cup Z)$. Then $\mathcal{D}_{Y,r}=\{Z\cup D\colon D\in\mathcal{D}^{\prime}\}$. Therefore, it suffices to find a vertex cover of $G^{\prime}$ of size $r-|Z|$, which can be implemented in FPT time parameterized by $r-|Z|\leq l$.
5.1.2 $t$-Hitting Set
Let $U$ be a finite set, $l,t\in\mathbb{Z}_{\geq 0}$, and $\mathcal{S}\subseteq 2^{U}$ be a set family such that each $S\in\mathcal{S}$ satisfies $|S|\leq t$. A subset $Z$ of $U$ is a hitting set of $\mathcal{S}$ if for each $S\in\mathcal{S}$, $S\cap Z\neq\emptyset$. Let $\mathcal{D}\subseteq 2^{U}$ be the family of all hitting sets of $\mathcal{S}$ of size at most $l$. [baste2019fpt] provided an FPT algorithm parameterized by $k,l,t$ for the max-min/max-sum diversification problems on $\mathcal{D}$.
We design an exact empty extension oracle on $\mathcal{D}$. Let $Y\subseteq 2^{U}$. The problem of finding a hitting set disjoint from $Y$ of size $r\in\mathbb{Z}_{\geq 0}$ is equivalent to finding a hitting set of size $r$ for the set family $\{S\setminus Y\colon S\in\mathcal{S}\}$. Therefore, the oracle can be designed using any FPT algorithm for the hitting set problem.
5.1.3 Feedback Vertex Set
Let $G=(V,E)$ be an undirected graph. A vertex subset $Z\subseteq V$ is a feedback vertex set of $G$ if the graph induced by $V\setminus Z$ is acyclic. Let $\mathcal{D}\subseteq 2^{V}$ be the set of all feedback vertex sets of $G$. [baste2019fpt] provided an FPT algorithm parameterized by $k,l$ for the max-min/max-sum diversification problems on $\mathcal{D}$.
We design an exact empty extension oracle on $\mathcal{D}$. Let $r\in\mathbb{Z}_{\geq 0}$ and $Y\subseteq V$. Since adding vertices to a feedback vertex set remains it to be a feedback vertex set, it suffices to find a feedback vertex set of size at most $r$ disjoint from $Y$. We use the compact representation of feedback vertex sets introduced by [guo2006compression] to achieve this. A family $\mathcal{C}$ of pairwise disjoint subsets of $V$ is said to represent a set $Z\subseteq V$ if for each $C\in\mathcal{C}$, $|Z\cap C|=1$. When all sets represented by $\mathcal{C}$ are feedback vertex sets of $G$, we call $\mathcal{C}$ a compact representation of feedback vertex sets. A list $\mathfrak{L}$ consisting of compact representations of feedback vertex sets, each of size at most $r$, is called a compact list of $r$-minimal feedback vertex sets if, for every inclusion-wise minimal feedback vertex set $Z$ of $G$, there exists $\mathcal{C}\in\mathfrak{L}$ that represents $Z$. The following is known.
Lemma 5.1 ([guo2006compression]).
There exists an algorithm that computes a compact list of $r$-minimal feedback vertex sets of size $2^{O(r)}poly(|V|)$ in $2^{O(r)}poly(|V|)$ time.
Our algorithm first computes a compact list $\mathfrak{L}$ of $r$-minimal feedback vertex sets. Then, for each $\mathcal{C}\in\mathfrak{L}$, the algorithm determines whether $\mathcal{C}$ represents a feedback vertex set disjoint from $Y$. This can be acheived by checking the non-emptiness of $C\setminus Y$ for each $C\in\mathcal{C}$. Since every minimal feedback vertex set is represented by some $\mathcal{C}\in\mathfrak{L}$, this procedure gives the desired oracle.
5.1.4 Matroid Intersection
Let $U$ be a finite set, $l\in\mathbb{Z}_{\geq 0}$, and $\mathcal{M}_{1}:=(U,\mathcal{I}_{1})$, $\mathcal{M}_{2}:=(U,\mathcal{I}_{2})$ be two matroids. A subset $Z$ of $U$ is a common independent set of $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ if $Z\in\mathcal{I}_{1}\cap\mathcal{I}_{2}$. Let $\mathcal{D}\subseteq 2^{U}$ be the set of all common independent sets of $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$. [fomin2024diversecollection] provided an FPT algorithm parameterized by $k,l$ for the max-min diversification problems on $\mathcal{D}$.
We design an exact empty extension oracle on $\mathcal{D}$. Let $Y\subseteq 2^{U}$. The problem of finding a common independent set of size $r\in\mathbb{Z}_{\geq 0}$ disjoint from $Y$ is equivalent to finding a common independent set in matroids $\mathcal{M}_{1}^{\prime}:=(U\setminus Y,\{I\colon I\cap Y=\emptyset\land I\in\mathcal{I}_{1}\})$ and $\mathcal{M}_{2}^{\prime}:=(U\setminus Y,\{I\colon I\cap Y=\emptyset\land I\in\mathcal{I}_{2}\})$. The oracle can thus be designed using any polynomial-time algorithm for the matroid intersection problem. Furthermore, the following result [marx2009parameterized] allows extending this result to when $\mathcal{D}$ is a family of common independent sets of $t\in\mathbb{Z}_{\geq 1}$ represented linear matroids.
Lemma 5.2 ([marx2009parameterized]).
There is an FPT algorithm parameterized by $t$ and $l$ for finding a common independent set of size $l$ of $t$ given represented linear matroids.
5.1.5 Reduction from Loss-Less Kernel-Based Framework
[baste2022diversity] constructed a framework using loss-less kernels [carbonnel2016propagation, carbonnel2017kernelization] for the max-sum diversification problem parameterized by $k,l$, which also straightforwardly applies to the max-min diversification problem. Here, we introduce a rephrased version of the loss-less kernel using our terminology. Let $U$ be a finite set and $\mathcal{D}\subseteq 2^{U}$ be a family such that $|D|\leq l$ holds for all $D\in\mathcal{D}$. A tuple $(\mathcal{D}^{\prime},P,Q)$ consisting of two disjoint subsets $P,Q\subseteq U$ and $\mathcal{D}^{\prime}\subseteq 2^{U\setminus(P\cup Q)}$ is a loss-less kernel of $\mathcal{D}$ if $\mathcal{D}=\{D^{\prime}\cup P\cup Q^{\prime}\colon D^{\prime}\in\mathcal{D}^{\prime}\land Q^{\prime}\subseteq Q\land|D\cup P\cup Q^{\prime}|\leq l\}$ and $|\mathcal{D}^{\prime}|$ is bounded by a constant that depends only on $l$.
Here, we observe that our result generalizes this framework. We design an exact empty extension oracle using the loss-less kernel. Let $r\in\mathbb{Z}_{\geq 0}$ and $Y\subseteq U$. Let $(\mathcal{D}^{\prime},P,Q)$ be the loss-less kernel of $\mathcal{D}$. The oracle exhaustively searches for $D^{\prime}\in\mathcal{D}^{\prime}$ and determines whether $Y\cap(D^{\prime}\cup P)=\emptyset$. If so, and if $|D^{\prime}\cup P|\leq r\leq|D^{\prime}\cup P\cup(Q\setminus Y)|$, a desired set is found. If no desired set is found for any $D^{\prime}\in\mathcal{D}^{\prime}$, the oracle returns $\bot$.
5.1.6 Reduction from Color-Coding-Based Framework
[hanaka2021finding] constructed a color-coding [alon1995color] based framework to provide FPT algorithms for the max-min diversification problems parameterized by $k,l$. Specifically, they used the following oracle to solve the max-min diversification problem.
Colorful Set Oracle on $\mathcal{D}$: Let $U$ be a finite set, $s,t\in\mathbb{Z}_{\geq 1}$ with $s\leq t$, $c\in\{1,\dots,t\}^{U}$, and $\mathcal{D}\subseteq 2^{U}$. It determines whether there exists a set $D\in\mathcal{D}$ of size $s$ such that $\{c_{e}\colon e\in D\}=\{1,\dots,s\}$, and if so, finds one.
They proved that there is an oracle algorithm for the max-min diversification problem on $\mathcal{D}$ using the colorful set oracle on $\mathcal{D}$, where the number of oracle calls and the time complexity are both FPT parameterized by $k,l$. They applied this framework on several domains $\mathcal{D}$, including the family of paths, interval schedulings, and matchings of size $l$.
Here, we observe that our framework generalizes theirs. It suffices to design an exact empty extension oracle on $\mathcal{D}$ using the colorful set oracle on $\mathcal{D}$. Let $U$ be a finite set, $r\in\mathbb{Z}_{\geq 0}$, and $Y\subseteq U$. Let $s=r,t=r+1$, and for each $e\not\in Y$, take $c_{e}\in\{1,\dots,r\}$ independently and uniformly at random. For $e\in Y$, let $c_{e}=r+1$. If $D\in\mathcal{D}$ satisfies $|D|=r$ and $D\cap Y=\emptyset$, $\{c_{e}\colon e\in D\}=\{1,\dots,s\}$ holds with probability $\frac{1}{r!}$. Thus, by calling the colorful set oracle on $\mathcal{D}$ $O(r!)$ times, we can implement the exact empty extension oracle that works correctly with sufficiently high probability. Furthermore, we can derandomize this algorithm using a perfect hash family [alon1995color].
5.2 $k,d$ as Parameters
Here we apply Theorem 1.2 to specific domains $\mathcal{D}$. For each domain $\mathcal{D}$, we design the $(-1,1)$-optimization oracle on $\mathcal{D}$ and the exact extension oracle on $\mathcal{D}$.
5.2.1 Matroid Base and Branching
Let $\mathcal{M}=(U,\mathcal{I})$ be a matroid and $\mathcal{D}$ be the family of bases of $\mathcal{M}$. [fomin2024diversecollection] provided an FPT algorithm parameterized by $k,d$ for the max-min diversification problem on $\mathcal{D}$. We design the $(-1,1)$-optimization oracle and the exact extension oracle on $\mathcal{D}$ to obtain an FPT algorithm for diversification and clustering problems on $\mathcal{D}$.
The $(-1,1)$-optimization oracle on $\mathcal{D}$ is just a weighted maximization problem on matroid bases and can be implemented in polynomial time using a greedy algorithm. We provide a polynomial-time algorithm for the exact extension oracle on $\mathcal{D}$. Let $r\in\mathbb{Z}_{\geq 0}$, $X,Y\subseteq U$, and $C\in\mathcal{D}$. If $r$ is odd, there is clearly no solution, so we assume $r$ is even. The algorithm first computes a set $D_{\min}\in\mathcal{D}$ (resp. $D_{\max}\in\mathcal{D}$) that minimizes (resp. maximizes) $|D\triangle C|$ among $D\in\mathcal{D}$ that contain $X$ and disjoint from $Y$. This can be formulated as a weighted maximization problem on a matroid and implemented using a greedy algorithm. If $r<|D_{\min}\triangle C|$ or $|D_{\max}\triangle C|<r$, there is no solution. Otherwise, we can find a solution using strong exchange property of $\mathcal{D}$. Here, a set family $\mathcal{Z}$ satisfies the strong exchange property if for any distinct $Z_{1},Z_{2}\in\mathcal{Z}$ and any $e_{1}\in Z_{1}\setminus Z_{2}$, there exists $e_{2}\in Z_{2}\setminus Z_{1}$ such that $Z_{1}\setminus\{e_{1}\}\cup\{e_{2}\}\in\mathcal{Z}$. It is well-known that the base family of a matroid satisfies the strong exchange property. Starting from $D_{1}=D_{\min}$ and $D_{2}=D_{\max}$, the algorithm repeatedly applies the strong exchange property on $D_{1}$ and $D_{2}$ and replaces $D_{1}$ by the obtained solution. Then, $|D_{1}\triangle C|=r$ eventually holds after at most $\frac{|D_{\min}\triangle D_{\max}|}{2}$ iterations.
We can easily see that this algorithm actually works under a weaker assumption, weak exchange property of $\mathcal{D}$. Here, $\mathcal{Z}$ satisfies the weak exchange property if for any distinct $Z_{1},Z_{2}\in\mathcal{Z}$, there exist elements $e_{1}\in Z_{1}\setminus Z_{2}$ and $e_{2}\in Z_{2}\setminus Z_{1}$ such that $Z_{1}\setminus\{e_{1}\}\cup\{e_{2}\}\in\mathcal{Z}$. Particularly, consider the case that $\mathcal{D}$ is the family of $r$-branching in a graph. Here, for a directed graph $G=(V,E)$ and $r\in V$, an edge subset $T\subseteq E$ is an $r$-branching if $|T|=|V|-1$ and for all vertices $v\in V$, there is an $r-v$ path consisting only of edges in $T$. The $(-1,1)$-optimization oracle on $\mathcal{D}$ can be implemented using any polynomial-time algorithm for finding a maximum weight $r$-branching. Since it is well-known that $\mathcal{D}$ satisfies weak exchange property [ito2023reconfiguring], we can design an exact extension oracle that works in polynomial time. Furthermore, when $\mathcal{D}$ is the family of branchings in a graph $G$, that is, the family of edge subsets forming an $r$-branching for some $r\in V$, $\mathcal{D}$ can be written as the union of the families of $r$-branchings for each $r\in V$. Hence, by the observation at the beginning of this section, the same result holds for $\mathcal{D}$ as well.
Given a directed graph $G=(V,E)$ and $s,t\in V$, the $d$-distinct branchings problem is to determine whether there exist an $s$-branching $B_{1}$ and a $t$-in-branching $B_{2}$ of $G$ such that $|B_{1}\triangle B_{2}|\geq d$, where an edge subset is an $r$-in-branching if it is an $r$-branching in the graph obtained by reversing the direction of every edge in $G$. [bang2016parameterized] provided an FPT algorithm for this problem when $G$ is strongly connected. [gutin2018k] extended their results to general graphs, and [bang2021k] obtained a polynomial kernel and improved the time complexity. By the discussion in Section 2.1, we can obtain another FPT algorithm for this problem using $d$-limited $k$-max-distance sparsifiers. Specifically, the algorithm computes the $d$-limited $1$-max-distance sparsifiers $\mathcal{K}_{1}$ and $\mathcal{K}_{2}$ of a family of $s$-branchings and $t$-in-branchings, respectively, with respect to $2^{U}$, and then performs an exhaustive search over pairs $(B_{1},B_{2})\in\mathcal{K}_{1}\times\mathcal{K}_{2}$. We can naturally extend this algorithm to the case where we choose multiple branchings and in-branchings. Specifically, we obtain an FPT algorithm parameterized by $k$ and $d$ for the following problem. Let $k\in\mathbb{Z}_{\geq 1}$, $d\in\mathbb{Z}_{\geq 0}$, $k^{\prime}$ be an integer with $0\leq k^{\prime}\leq k$, and $s_{1},\dots,s_{k}\in V$. Let $\mathcal{D}_{1},\dots,\mathcal{D}_{k^{\prime}}$ be the family of $s_{1},\dots,s_{k^{\prime}}$-branchings, respectively, and $\mathcal{D}_{k^{\prime}+1},\dots,\mathcal{D}_{k}$ be the family of $s_{k^{\prime}+1},\dots,s_{k}$-in-branchings, respectively. The problem asks whether there is a $k$-tuple $(B_{1},\dots,B_{k})\in\mathcal{D}_{1}\times\dots\times\mathcal{D}_{k}$ such that $\min_{1\leq i<j\leq k}|D_{i}\triangle D_{j}|\geq d$ (or $\sum_{1\leq i<j\leq k}|D_{i}\triangle D_{j}|\geq d$).
5.2.2 Matching
Let $G=(V,E)$ be an undirected graph. An edge subset $M\subseteq E$ is a matching of $G$ if no two edges in $M$ share a vertex. A matching $M$ of $G$ is perfect if $2|M|=|V|$. Let $l\in\left\{0,\dots,\left\lfloor{\frac{|V|}{2}}\right\rfloor\right\}$ and $\mathcal{D}$ be the family of matchings in $G$ consists of $l$ edges. [fomin2024diversepair] provided FPT algorithms parameterized by $d$ for the max-min diversification problem on $\mathcal{D}$ for the case $k=2$ and $l$ is the size of the maximum matching. [fomin2024diversecollection] provided the same result parameterized by $k,d$ for general $k$ for the case $2l=|V|$. Let $d\in\mathbb{Z}_{\geq 0}$ and $k\in\mathbb{Z}_{\geq 1}$. In this section, we design the $(-1,1)$-optimization oracle and the exact extension oracle on $\mathcal{D}$. The following lemma is useful.
Lemma 5.3.
Let $W$ a vertex set of size $|V|-2l$ disjoint from $V$. Let $F:=\{(v,w)\colon v\in V,w\in W\}$ and $G^{l}:=(V\dot{\cup}W,E\dot{\cup}F)$. Then, if $M^{l}\subseteq E\dot{\cup}F$ is a perfect matching in $G^{l}$, $M^{l}\cap E$ is a matching of size $l$ in $G$. Conversely, if there exists a matching of size $l$ in $G$, there exists a perfect matching in $G^{l}$.
Proof.
Let $M^{l}$ be a perfect matching in $G^{l}$. Since $W$ is an independent set, $M^{l}$ consists of $|V|-2l$ edges between $V$ and $W$ and $l$ edges within $V$, which means $M^{l}\cap E$ is a matching of size $l$ in $G$. Conversely, a matching of size $l$ in $G$ can be naturally extended to a perfect matching in $G^{l}$. ∎
We call $G^{l}$ the $l$-expanded graph of $G$. We now design a $(-1,1)$-optimization oracle on $\mathcal{D}$ that runs in polynomial time. Let $w\in\{-1,1\}^{E}$. The algorithm first constructs the $l$-expanded graph $G^{l}:=(V\dot{\cup}W,E\dot{\cup}F)$ of $G$. Define the edge weights $\widehat{w}\in\{-1,0,1\}^{E\dot{\cup}F}$ by $\widehat{w}_{e}=w_{e}$ for $e\in E$ and $\widehat{w}_{e}=0$ for $e\in F$. Then, the algorithm finds the maximum weight perfect matching $M^{l}$ with respect to $\widehat{w}$. From Lemma 5.3, $M^{l}\cap E$ is the desired output for the $(-1,1)$-optimization oracle on $\mathcal{D}$.
We now design an exact extension oracle on $\mathcal{D}$ that runs in polynomial time. Let $r\in\mathbb{Z}_{\geq 0}$, $X,Y\subseteq E$, and $C\in\mathcal{D}$. Let $G^{\prime}:=(V^{\prime}:=V\setminus V[X],E^{\prime}:=E\setminus(X\cup Y))$, where $V[X]$ denotes the set of endpoints of edges in $X$. Color each edge $e\in E^{\prime}$ red if $e\in C$ and blue if $e\notin C$. The problem reduces to determining whether there exists a matching of size $l-|X|$ in $G^{\prime}$ that contains exactly $r-|X\setminus C|$ blue edges. Let $G^{\prime\prime}=(V^{\prime}\cup W^{\prime},E^{\prime}\cup F^{\prime})$ be the $(l-|X|)$-expanded graph of $G^{\prime}$, and color the edges in $F^{\prime}$ red. The problem reduces to determining whether there exists a perfect matching in $G^{\prime\prime}$ containing exactly $r-|X\setminus C|$ blue edges. This problem is known as the exact matching problem [mulmuley1987matching] and admits a randomized polynomial-time algorithm.
5.2.3 Minimum Edge Flow
Let $G=(V,E)$ be a graph. For simplicity, we consider only the case when $G$ is directed; the results in this part naturally extend to undirected graphs. Let $s,t\in V$ and $b\in\mathbb{Z}_{\geq 0}$. An edge subset $F\subseteq E$ is a $s,t$-flow of amount $b$ if $F$ can be written as the union of the edge sets of $b$ edge-disjoint $s,t$-paths. Let $\mathcal{D}$ be the family of $s,t$-flows of amount $b$ with the minimum number of edges. In this section, we design the $(-1,1)$-optimization oracle on $\mathcal{D}$ and the exact extension oracle on $\mathcal{D}$.
We provide a $(-1,1)$-optimization oracle on $\mathcal{D}$ that runs in polynomial time. Let $w\in\{-1,1\}^{E}$. Define edge weights $\widehat{w}\in\mathbb{Z}^{E}$ by $\widehat{w}_{e}:=2|E|-w_{e}$. The algorithm finds the minimum weight $s,t$-flow $F$ of amount $b$ with respect to $\widehat{w}$. The correctness of the algorithm is shown as follows.
Lemma 5.4.
$F$ is the desired output for the $(-1,1)$-optimization oracle on $\mathcal{D}$.
Proof.
It is sufficient to state that $F$ is an $s,t$-flow of amount $b$ with the minimum number of edges. Assume the contrary and let $F^{\prime}$ be an $s,t$-flow of amount $b$ with the minimum number of edges. Then, we have $\widehat{w}(F^{\prime})\leq 2|E||F^{\prime}|+|F^{\prime}|\leq 2|E||F|-(2|E|-|F^{\prime}|)<2|E||F|-|F|\leq\widehat{w}(F)$, contradicting the minimality of $\widehat{w}(F)$. ∎
We provide an exact extension oracle on $\mathcal{D}$ that runs in FPT time. Let $r\in\mathbb{Z}_{\geq 0}$, $X,Y\subseteq E$, and $F\in\mathcal{D}$. Let $G^{\prime}=(V,E^{\prime})$ be the graph obtained from $G$ by reversing the direction of edges in $F$. For an edge set $Z\subseteq E^{\prime}$ and a vertex $v\in V$, let $\delta^{+}_{Z}(v)$ denote the set of edges in $Z$ leaving $v$ and $\delta^{-}_{Z}(v)$ denote the set of edges in $Z$ entering $v$. An edge set $Z$ of a directed graph is Eulerian if for all $v\in V$, $|\delta^{+}_{Z}(v)|=|\delta^{-}_{Z}(v)|$. Let $\mathcal{D}^{*}$ be the family of Eulerian edge sets $D^{*}$ such that $|D^{*}\cap F|=|D^{*}\setminus F|$. Then $\mathcal{D}=\{F\triangle D^{*}\colon D^{*}\in\mathcal{D}^{*}\}$. Thus, to design the exact extension oracle on $\mathcal{D}$, it suffices to determine whether there exists an Eulerian edge subset $D^{*}\subseteq E^{\prime}$ such that $|D^{*}\cap F|=|D^{*}\setminus F|=\frac{r}{2}$, $X^{\prime}\subseteq Z$, and $Y^{\prime}\cap Z=\emptyset$, where $X^{\prime}:=(X\setminus F)\cup(Y\cap F)$ and $Y^{\prime}:=(X\cap F)\cup(Y\setminus F)$. Using the fact that Eulerian edge sets can be decomposed into edge-disjoint cycles, we solve this problem via color-coding. For each $e\in E^{\prime}$, sample $c_{e}\in\{1,\dots,r\}$ independently and uniformly at random. If there exists a $D^{*}\in\mathcal{D}^{*}$ meeting the conditions, there is a $\frac{1}{r!}$ probability that $\{c_{e}\colon e\in D^{*}\}=\{1,\dots,r\}$. Define $\mathsf{DP}[u][v][C][p][q]$ by $Z$ if there exists a $Z\subseteq E^{\prime}\setminus Y^{\prime}$ that satisfies the following conditions, and $\bot$ otherwise.
  • $|\delta^{+}_{Z}(w)|-|\delta^{-}_{Z}(w)|$ is equal to $1$ if $w=u\neq v$, $-1$ if $w=v\neq u$, and 00 otherwise,
  • $|Z|=|C|$ and $\{c_{e}\colon e\in Z\}=C$,
  • $|Z\cap X^{\prime}|=p$, and
  • $|Z\cap F|=q$.
If there exists $v\in V$ such that $\mathsf{DP}[v][v][\{1,\dots,r\}][|X^{\prime}|][r/2]\neq\bot$, then this is the desired output for the oracle. The algorithm initializes the table with
$\mathsf{DP}[u][v][C][p][q]:=\begin{cases}\emptyset&(u=v,C=\emptyset,p=q=0)\\ \bot&(\text{otherwise})\end{cases}$
and updates the table by the following rules.
  • If $\mathsf{DP}[u][v][C][p][q]\neq\bot$, set $\mathsf{DP}[u][w][C\cup\{c_{e}\}][p^{\prime}][q^{\prime}]:=\mathsf{DP}[u][v][C][p][q]\cup\{e\}$ for each $w\in V$ such that there exists an edge $e:=(v,w)\in E^{\prime}$ with $c_{e}\not\in C$, where $p^{\prime}$ equals $p+1$ if $e\in X^{\prime}$ and $p$ otherwise, and $q^{\prime}$ equals $q+1$ if $e\in F$ and $q$ otherwise.
  • If $\mathsf{DP}[v][v][C][p][q]\neq\bot$, set $\mathsf{DP}[w][w][C][p][q]:=\mathsf{DP}[v][v][C][p][q]$ for each $w\in V$.
This algorithm runs in FPT time parameterized by $r$.
5.2.4 Minimum Edge Steiner Tree
Let $G=(V,E)$ be an undirected graph and $T\subseteq V$. An edge subset $F\subseteq E$ is called a Steiner tree with a terminal set $T$ if the graph $\left(V_{F}:=\bigcup_{e\in F}e,F\right)$ is connected and $T\subseteq V_{F}$. Let $\mathcal{D}$ be the family of all Steiner trees with the minimum number of edges. In this section, we design the $(-1,1)$-optimization oracle on $\mathcal{D}$ and the exact extension oracle on $\mathcal{D}$. [funayama2024parameterized] provided an FPT algorithm for the max-min diversification problem on the domain of shortest paths parameterized by $k$ and $d$. Their problem for undirected graphs corresponds to the case of $|T|=2$.
We provide a $(-1,1)$-optimization oracle on $\mathcal{D}$ that runs in FPT time parameterized by $|T|$. This is almost identical to the textbook algorithm [cygan2015parameterized]. Let $w\in\{-1,1\}^{E}$. For each $v\in V$ and $S\subseteq T$, let $\mathsf{OPT}[v][S]$ be the minimum number of edges in a Steiner tree with terminal set $S\cup\{v\}$, and $\mathsf{DP}[v][S]$ be the one with the maximum weight among such Steiner trees. What we seek is $\mathsf{DP}[v][T]$ for any $v\in T$. This dynamic programming algorithm initializes
$\mathsf{OPT}[v][S]:=\begin{cases}0&(v\in T,S=\{v\})\\ \infty&(\text{otherwise})\end{cases},\quad\quad\mathsf{DP}[v][S]:=\begin{cases}\emptyset&(v\in T,S=\{v\})\\ \bot&(\text{otherwise})\end{cases}$
and updates using
$\mathsf{OPT}[v][S]$ $:=\min\left(\min_{u\in V\colon(u,v)\in E}\mathsf{OPT}[u][S]+1,\min_{\emptyset\neq S^{\prime}\subsetneq S}\left(\mathsf{OPT}[v][S^{\prime}]+\mathsf{OPT}[v][S\setminus S^{\prime}]\right)\right),$
to compute $\mathsf{OPT}[\cdot][\cdot]$ and sets $\mathsf{DP}[v][S]$ to be the $\bot$ set with the maximum weight among the following sets, if any; otherwise, $\bot$.
  • $\mathsf{DP}[u][S]\cup\{(u,v)\}$ for each $u\in V$ such that $(u,v)\in E$ and $\mathsf{OPT}[u][S]+1=\mathsf{OPT}[v][S]$.
  • $\mathsf{DP}[v][S^{\prime}]\cup\mathsf{DP}[v][S\setminus S^{\prime}]$ for each $\emptyset\neq S^{\prime}\subsetneq S$ such that $\mathsf{OPT}[v][S^{\prime}]+\mathsf{OPT}[v][S\setminus S^{\prime}]=\mathsf{OPT}[v][S]$.
Here, for simplicity, we set $\bot\cup Z=\bot$ for any set $Z$. This runs in FPT time parameterized by $|T|$.
We provide an exact extension oracle on $\mathcal{D}$ that runs in FPT time parameterized by $|T|$. This is also almost identical to the same textbook algorithm. Let $r\in\mathbb{Z}_{\geq 0}$, $X,Y\subseteq E$, and $F\in\mathcal{D}$. For each $v\in V$, $S\subseteq T$, $p\in\{0,\dots,|X|\}$, and $q\in\{0,\dots,r/2\}$, define $\mathsf{EX}[v][S][p][q]$ as the Steiner tree with $\mathsf{OPT}[v][S]$ edges with terminal set $S\cup\{v\}$, disjoint from $Y$, including $p$ edges in $X$, and including $q$ edges not in $F$, if such Steiner tree exists; otherwise, it is $\bot$. The solution required for the oracle is $\mathsf{EX}[v][T][|X|][r/2]$ for any $v\in T$. This dynamic programming algorithm initializes
$\mathsf{EX}[v][S][p][q]:=\begin{cases}\emptyset&(v\in T,S=\{v\},p=q=0)\\ \bot&(\text{otherwise})\end{cases}$
and sets $\mathsf{EX}[v][S][p][q]$ to be the $\bot$ sets among the following if there exists any; otherwise, $\bot$.
  • $\mathsf{EX}[u][S][p^{\prime}][q^{\prime}]\cup\{(u,v)\}$ for each $u\in V$ such that $(u,v)\in E$ and $\mathsf{OPT}[u][S]+1=\mathsf{OPT}[v][S]$. Here, $p^{\prime}$ equals $p-1$ if $(u,v)\in X$, and $p$ otherwise; $q^{\prime}$ equals $q-1$ if $(u,v)\not\in F$, and $q$ otherwise.
  • $\mathsf{EX}[v][S^{\prime}][p^{\prime}][q^{\prime}]\cup\mathsf{EX}[v][S\setminus S^{\prime}][p-p^{\prime}][q-q^{\prime}]$ for each $\emptyset\neq S^{\prime}\subsetneq S$ with $\mathsf{OPT}[v][S^{\prime}]+\mathsf{OPT}[v][S\setminus S^{\prime}]=\mathsf{OPT}[v][S]$, $p^{\prime}\in\{0,\dots,p\}$, and $q^{\prime}\in\{0,\dots,q\}$.
This dynamic programming algorithm runs in FPT time parameterized by $|T|$.
5.2.5 Vertex Set of Minimum $s,t$-Cut
Let $G=(V,E)$ be a directed or undirected graph and $s,t\in V$. A vertex subset $C\subseteq V$ is an $s,t$-cut if $s\in C$ and $t\not\in C$. The cost of the $s,t$-cut $C$ is $|\delta(C)|:=|\{e\in E\colon|e\cap C|=1\}|$. Let $\mathcal{D}\subseteq 2^{V}$ be the family of minimum cost $s,t$-cuts. Let $k\in\mathbb{Z}_{\geq 1}$ and $d\in\mathbb{Z}_{\geq 0}$. In this section, we construct a $d$-limited $k$-max-distance sparsifier of $\mathcal{D}$.
We provide an $(-1,1)$-optimization oracle that runs in polynomial time. Let $w\in\{-1,1\}^{V}$. We construct the edge set $\widehat{E}$ consisting of
  • for each $e\in E$, an edge $e$ with weight $12|V|+$ 1,
  • for each $v\in V$ with $w_{v}=1$, an edge $(s,v)$ with weight $1$, and
  • for each $v\in V$ with $w_{v}=-1$, an edge $(v,t)$ with weight $1$.
Then, solve the minimum weight $s,t$-cut problem on the graph $(V,\widehat{E})$. We can easily see that the $s,t$-cut returned by this procedure is indeed the solution required by the oracle by the similar argument as the proof of Lemma 5.4.
Let $C\in\mathcal{D}$, $r\in\mathbb{Z}_{\geq 0}$ and $X,Y\subseteq U$. Unfortunately, we do not have an exact extension oracle itself; however, we can construct an FPT algorithm that either outputs a trivial $d$-limited $k$-max-distance sparsifier or behaves as an exact extension oracle. Given that our goal in this section is to compute the $d$-limited $k$-max-distance sparsifier, this is sufficient. Furthermore, if this procedure returns a trivial $d$-limited $k$-max-distance sparsifier, it implies there is no solution to the $k$-center or $k$-sum-of-radii clustering problems. Therefore, algorithms for clustering problems can also be obtained without any issues.
A finite set equipped with a binary relation $(W,\preceq)$ is a poset if $\preceq$ is reflexive, antisymmetric, and transitive. For $v\in W$, define $I^{+}(v)=\{u\in W\colon u\succeq v\}$ and $I^{-}(v)=\{u\in W\colon u\preceq v\}$. A subset $I\subseteq W$ is called an ideal if for any $w\in I$, $I^{-}(w)\subseteq I$. The following is well-known.
Lemma 5.5 ([birkhoff1937rings, BergMS23]).
There is a poset $\mathcal{P}=(W,\preceq)$ and a list of nonempty disjoint subsets $\{A_{w}\}_{w\in W}$ of $U$ such that $\mathcal{D}=\bigcup_{\text{$I$ is ideal of $\mathcal{P}$}}\left\{\bigcup_{w\in I}A_{w}\right\}$. Such a poset $\mathcal{P}$ can be computed in polynomial time.
We take the poset $\mathcal{P}=(W,\preceq)$ and $\{A_{w}\}_{w\in W}$ in Lemma 5.5 and the ideal $I_{C}$ corresponding to $C$. Let $p:=(4d+2)^{2}2^{k-1}$ be the upper bound of $r$ among all calls of the exact extension oracle in our framework. Let $W^{+}_{C}:=\left\{w\in W\setminus I_{C}\colon|I^{-}(w)\setminus I_{C}|\leq p\right\}$ and $W^{-}_{C}:=\left\{w\in I_{C}\colon|I^{+}(w)\cap I_{C}|\leq p\right\}$. We have the following.
Lemma 5.6.
Let $D\in\mathcal{D}$ and assume $|D\triangle C|\leq p$. Let $I_{D}$ be the ideal of $\mathcal{P}$ corresponding to $D$. Then, $I_{C}\setminus W^{-}_{C}\subseteq I_{D}\subseteq I_{C}\cup W^{+}_{C}$.
Proof.
Assume there is an element $w$ in $I_{D}\setminus(I_{C}\cup W^{+}_{C})$. From the definition of $W^{+}_{C}$, we have $|I^{-}(w)\setminus I_{C}|>p$. Therefore, $|D\triangle C|\geq\sum_{w\in I_{D}\setminus I_{C}}|A_{w}|\geq|I_{D}\setminus I_{C}|\geq|I^{-}(w)\setminus I_{C}|>p$. Similarly, assume there is an element $w$ in $(I_{C}\setminus W^{-}_{C})\setminus I_{D}$. From the definition of $W^{-}_{C}$, we have $|I^{+}(w)\cap I_{C}|>p$. Therefore, $|D\triangle C|\geq\sum_{w\in I_{C}\setminus I_{D}}|A_{w}|\geq|I_{C}\setminus I_{D}|\geq|I_{C}\cap I^{+}(w)|>p$. ∎
Since the size of $W^{+}_{C}$ or $W^{-}_{C}$ can be large in general, Lemma 5.6 does not directly lead to an FPT algorithm for the exact extension oracle. However, we can claim the following.
Lemma 5.7.
Assume $|W^{+}_{C}|\geq 2kd$. Then, a $d$-limited $k$-max-distance sparsifier of $\mathcal{D}$ can be computed in polynomial time. The same holds for $|W^{-}_{C}|$.
Proof.
Order elements of $W^{+}_{C}$ as $w_{1},\dots,w_{|W^{+}_{C}|}$ so that $w_{j}\preceq w_{j^{\prime}}$ implies $j\leq j^{\prime}$. For each $j\in\{0,\dots,|W^{+}_{C}|\}$, $I_{C}\cup\{w_{1},\dots,w_{j}\}$ is an ideal of $\mathcal{P}$. For $i\in\{0,\dots,k\}$, let $C_{i}:=C\cup\bigcup_{j\in\{1,\dots,2id\}}A_{w_{j}}$. Then, $\{C_{0},\dots,C_{k}\}$ is a $d$-limited $k$-max-distance sparsifier by Lemma 4.1. The same argument applies to $W^{-}_{C}$. ∎
Thus, we can assume $|W^{+}_{C}|,|W^{-}_{C}|<2kd$. Then, Lemma 5.6 implies that there is at most $2^{4kd}$ sets $D\in\mathcal{D}$ with $|D\triangle C|\leq p$, and thus the exact extension oracle can be designed by a brute-force search.
5.2.6 Vertex Set of Edge Bipartization
Let $G=(V,E)$ be an undirected graph. For $D\subseteq V$, the edge bipartization set of $D$ is defined by $\beta(D):=E\setminus\delta(D)$. Let $s:=\min_{D\subseteq U}|\beta(D)|$ and $\mathcal{D}$ be the family of all $D\subseteq V$ with $|\beta(D)|=s$. In this section, we design the $(-1,1)$-optimization oracle on $\mathcal{D}$ and the exact extension oracle on $\mathcal{D}$.
Take any $D\in\mathcal{D}$. We use the technique in [guo2006compression] for designing iterative compression algorithms for the edge bipartition problem to characterize the domain $\mathcal{D}$ using $D$. For $D_{1},D_{2}\subseteq V$, we denote $E(D_{1},D_{2}):=\{(u,v)\in E\colon u\in D_{1}\land v\in D_{2}\}$. We begin by the following.
Lemma 5.8.
Let $D_{1},D_{2}\subseteq V$. Then, $\beta(D_{1})\triangle\beta(D_{2})=\delta(D_{1}\triangle D_{2})$.
Proof.
Then, we have $\beta(D_{1})\setminus\beta(D_{2})=E(D_{1}\setminus D_{2},D_{1}\cap D_{2})\cup E(D_{2}\setminus D_{1},V\setminus(D_{1}\cup D_{2}))$. Similarly, $\beta(D_{2})\setminus\beta(D_{1})=E(D_{2}\setminus D_{1},D_{1}\cap D_{2})\cup E(D_{1}\setminus D_{2},V\setminus(D_{1}\cup D_{2}))$. Thus, $\beta(D_{1})\triangle\beta(D_{2})=E(D_{1}\triangle D_{2},D_{1}\cap D_{2})\cup E(D_{1}\triangle D_{2},V\setminus(D_{1}\cup D_{2}))=E(D_{1}\triangle D_{2},V\setminus(D_{1}\triangle D_{2}))=\delta(D_{1}\triangle D_{2})$. ∎
For $A,B\subseteq U$ and $t\in\mathbb{Z}_{\geq 0}$, define $\mathrm{MinCut}(A,B,t)$ as the family of all sets $C\subseteq V$ with $A\subseteq C$, $B\cap C=\emptyset$, and $|\delta(C)|=t$ if $t$ is the minimum possible value of $|\delta(C)|$; otherwise, define it as $\emptyset$. Furthermore, for $Z\subseteq E$, let $V[Z]:=\bigcup_{e\in Z}e$. We have the following.
Lemma 5.9.
For any $D\in\mathcal{D}$,
$\mathcal{D}=\bigcup_{T\subseteq V[\beta(D)]}\left\{D\triangle C\colon C\in\mathrm{MinCut}(T,V[\beta(D)]\setminus T,2|\beta(D)\cap\delta(T)|)\right\}.$
Proof.
Let $D^{\prime}\subseteq U$. Then, we have
$|\beta(D^{\prime})|=|\beta(D)|-2|\beta(D)\setminus\beta(D^{\prime})|+|\beta(D)\triangle\beta(D^{\prime})|=|\beta(D)|-2|\beta(D)\setminus\beta(D^{\prime})|+|\delta(D\triangle D^{\prime})|.$
Let $T\subseteq V[\beta(D)]$ and assume $V[\beta(D)]\cap(D\triangle D^{\prime})=T$. Then, we have
$\beta(D)\setminus\beta(D^{\prime})=E(D\setminus D^{\prime},D\cap D^{\prime})\cup E(D^{\prime}\setminus D,V\setminus(D\cup D^{\prime}))=\beta(D)\cap\delta(D\triangle D^{\prime})=\beta(D)\cap\delta(T),$
where the last equality is from the fact that
$F\cap\delta(Z)=F\cap((\delta(V[F]\cap Z)\setminus E(V[F]\cap Z,Z\setminus V[F]))\cup E(Z\setminus V[F],V\setminus Z))=F\cap\delta(V[F]\cap Z)$
holds for all $F\subseteq E$ and $Z\subseteq V$. By the minimality of $|\beta(D)|$, $|\beta(D^{\prime})|=|\beta(D)|=s$ holds if and only if $|\delta(D\triangle D^{\prime})|$ takes the minimum value $2|\beta(D)\cap\delta(T)|$. ∎
Lemma 5.9 implies that optimization problems on $\mathcal{D}$ can be reduced to solving the corresponding optimization problems on the domain $\mathrm{MinCut}(T,V[\beta(D)]\setminus T,t)$ for each of $2^{|\beta(D)|}\leq 4^{s}$ candidates of $T\subseteq\beta(D)$, where $t=2|\beta(D)\cap\delta(T)|\leq 2s$.
We provide the $(-1,1)$-optimization oracle on $\mathcal{D}$ that runs in FPT time parameterized by $s$. Let $w\in\{-1,1\}^{V}$. From Lemma 5.9, given two disjoint vertex subsets $A,B\subseteq V$ and $t\leq 2s$, it is sufficient to find a set that maximizes $w(C)$ among $C\in\mathrm{MinCut}(A,B,t)$. The algorithm constructs a graph $(\widehat{V},\widehat{E})$, where $\widehat{V}=V\dot{\cup}\{a,b\}$ and $\widehat{E}$ consists of the following edges.
  • For each $e\in E$, an edge $e$ with weight $12|V|+$ 1.
  • For each $a^{\prime}\in A$, an edge $(a,a^{\prime})$ with weight $\infty$.
  • For each $b^{\prime}\in B$, an edge $(b^{\prime},b)$ with weight $\infty$.
  • For each $v\in V$ with $w_{v}=1$, an edge $(a,v)$ with weight $1$.
  • For each $v\in V$ with $w_{v}=-1$, an edge $(v,b)$ with weight $1$.
Then, the algorithm solve the minimum weight $a,b$-cut problem on the graph $(\widehat{V},\widehat{E})$. If the weight of the returned cut is in the range $[(2|V|+1)t-|V|,(2|V|+1)t+|V|]$, it maximizes $w(C)$ among $C\in\mathrm{MinCut}(A,B,t)$. Otherwise, $\mathrm{MinCut}(A,B,t)=\emptyset$.
Let $r\in\mathbb{Z}_{\geq 0}$, $X,Y\subseteq V$, and $D\in\mathcal{D}$. We provide an exact extension oracle on $\mathcal{D}$ that runs in FPT time parameterized by $s$ and $r$. From Lemma 5.9, given two disjoint subsets $A,B\subseteq V$ and $t\leq 2s$, it is sufficient to find a set in $C\in\mathrm{MinCut}(A,B,t)$ such that $|C|=r$, $X\subseteq D\triangle C$, and $Y\cap(D\triangle C)=\emptyset$. Without loss of generality, we can assume $\mathrm{MinCut}(A,B,t)\neq\emptyset$. The algorithm first computes $C\in\mathrm{MinCut}(A,B,t)$ that is inclusion-wise minimal. It is well-known that such a cut is unique; specifically, it is the cut corresponding to the ideal $\emptyset$ in Lemma 5.5. If $|C|>r$, there is no set $C$ satisfying the condition, and the oracle returns $\bot$. If $|C|=r$, the algorithm checks the conditions $X\subseteq D\triangle C$ and $Y\cap(D\triangle C)=\emptyset$, and if satisfied, outputs $C$; otherwise, it returns $\bot$. Assume $|C|<r$. Since $C$ is unique inclusion-wise minimal set in $\mathrm{MinCut}(A,B,t)$, for any $C^{\prime}\in\mathrm{MinCut}(A,B,t)\setminus\{C\}$, there exist an edge $e\in\delta(C)$ with $C^{\prime}\in\mathrm{MinCut}(A\cup e,B,t)$. For each edge $e\in\delta(C)$, the algorithm recurses on the instance where $A$ is replaced by $A\cup e$ to compute the desired set $C$. Since $|\delta(C)|\leq t\leq 2s$ and the recursion depth is at most $r$, this recursion is called at most $s^{O(r)}$ times.
5.2.7 Dynamic Programming
Let $U$ be a finite set, $G=(V,E)$ be a directed acyclic graph, and $q\in V^{U}$ be a labeling of vertices such that no path in $G$ passes through multiple vertices with the same label. For a path $P$ in $G$, let $q(P)\subseteq U$ be the set of labels of the vertices in $P$. Let $\mathcal{D}\subseteq 2^{U}$ be the family of subsets $D\subseteq U$ such that there exists a longest path $P$ in $G$ with $q(P)=D$. This formulation captures the solution domains of several typical dynamic programming problems. In particular, the solution domain of the interval scheduling problem, for which [hanaka2021finding] provided an FPT algorithm for the max-min diversification problem parameterized by $k$ and $l$, is represented as
  • $U=V$ is the set of intervals,
  • $E$ is the set of pairs of intervals $(u,v)$ such that the right end of $u$ is to the left of the left end of $v$, and
  • $q_{v}=v$ for all $v\in V$.
The solution domain of the longest common subsequence problem, for which [shida2024finding] provided an FPT algorithm for the max-min diversification problem parameterized by $k$ and $l$ for the case the number $m$ of input strings $\{S_{1},\dots,S_{m}\}$ is an absolute constant, is represented as
  • $U=\{1,\dots,l\}\times\Sigma$, where $\Sigma$ is the set of letters appearing in the input,
  • $V=\{p:=(p_{1},\dots,p_{m})\colon s_{p}:=S_{1,p_{1}}=\dots=S_{m,p_{m}}\}$,
  • $E$ is the set of pairs $(p,p^{\prime})\in V\times V$ such that $p_{i}<p^{\prime}_{i}$ for all $i\in\{1,\dots,m\}$, and
  • $q_{p}=(\mathrm{len}_{p},s_{p})$ for all $p\in V$, where $\mathrm{len}_{p}$ represents the maximum number of vertices of a path that ends at $p$.
Let $k\in\mathbb{Z}_{\geq 1}$ and $d\in\mathbb{Z}_{\geq 0}$. In this section, we design the $(-1,1)$-optimization oracle on $\mathcal{D}$ and the exact extension oracle on $\mathcal{D}$. Particularly, by constructing these oracles, we obtain an FPT algorithm parameterized by $k$ and $d$ for the max-min diversification problem on $\mathcal{D}$, which is a stronger result in terms of parameterization compared to the result in [hanaka2021finding] for the interval scheduling problem and the result in [shida2024finding] for the longest common subsequence problem.
We provide a $(-1,1)$-optimization oracle on $\mathcal{D}$ that runs in polynomial time. Let $w\in\{-1,1\}^{U}$. This can be achieved by dynamic programming where for each $v\in V$, $\mathsf{OPT}[v]$ denotes the maximum number of vertices in a path ending at $v$, and $\mathsf{DP}[v]$ denotes the maximum total weight of the labels on a path ending at $v$ with $\mathsf{OPT}[v]$ vertices.
We provide an exact extension oracle on $\mathcal{D}$ that runs in polynomial time. Let $r\in\mathbb{Z}_{\geq 0}$, $X,Y\subseteq U$, and $C\subseteq U$. This can be achieved by dynamic programming where for each $v\in V$, $p\in\{0,\dots,|X|\}$, $q\in\{0,\dots,r\}$, $\mathsf{EX}[v][p][q]$ denotes a path ending at $v$ with $\mathsf{OPT}[v]$ vertices that does not include any vertex with its label in $Y$, includes $p$ vertices with labels in $X$, and includes $q$ vertices with labels not belonging to $C$, if such a path exists; otherwise, it is $\bot$.